### v1

# CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis

Abstract
Depression is a pressing global public health issue, yet publicly available Chinese-language resources for risk detection remain scarce and are mostly limited to binary classification. To address this limitation, we release CNSocialDepress, a benchmark dataset for depression risk detection from Chinese social media posts. The dataset contains 44,178 texts from 233 users, within which psychological experts annotated 10,306 depression-related segments. CNSocialDepress provides binary risk labels together with structured multi-dimensional psychological attributes, enabling interpretable and fine-grained analysis of depressive signals. Experimental results demonstrate its utility across a wide range of NLP tasks, including structured psychological profiling and fine-tuning of large language models for depression detection. Comprehensive evaluations highlight the dataset’s effectiveness and practical value for depression risk identification and psychological analysis, thereby providing insights to mental health applications tailored for Chinese-speaking populations.
CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis
Jinyuan XU1,*, Tian LAN2,*, Xintao YU3, Xue HE3,4, Hezhi ZHANG5, Ying WANG6, Pierre Magistry1, Mathieu Valette1, Lei LI7,8†thanks: * These authors contributed equally to this work.thanks: † Corresponding Author. 1 Ertim Inalco, 2 Milkuya Studio, 3 Sorbonne Université, 4 IRD Lab 5 Faculty of Psychology, Peking University 6 Faculty of Psychology and Cognitive Science, Beijing Normal University 7 University of Washington, 8 VitaSight
1 Introduction
Depressive disorders are among the most common mental health conditions worldwide, characterized by persistent low mood or loss of interest in daily activities. According to a 2023 report [^1] https://www.who.int/news-room/fact-sheets/detail/depression by the World Health Organization (WHO), approximately 280 million people globally suffer from depression. In China, a 2024 survey by the Chinese Center for Disease Control and Prevention Wang et al. (2024a) estimates that around 95 million individuals are affected by depression, and out of the 280,000 suicides reported annually, 40% are linked to depressive disorders. Moreover, research Wang et al. (2024a) shows that depression strongly correlates with both suicidal and non-suicidal self-injurious deaths.
Motivated by the urgent need for the detection of depression, researchers increasingly apply machine learning (ML) and natural language processing (NLP) methods for automated analysis of depression risk Squires et al. (2023); Hasib et al. (2023); Aleem et al. (2022). While early efforts have shown promising results, they are heavily constrained by the limitations of existing datasets. Traditional depression detection studies often rely on clinical dataBittar et al. (2019); Fernandes et al. (2018), or on transcripts from medical or psychological interviews resources Shen et al. (2022); Li et al. (2022) that are are expensive to collect, limited in size and diversity, and may not reflect the informal, emotionally nuanced expressions typical of real-world online environments.
To overcome these limitations, recent research has shifted toward leveraging user-generated content on social media platforms [^2] https://github.com/bucuram/depression-datasets-nlp Harrigian et al. (2021), which provides rich linguistic signals and greater availability. Such data are also more easily anonymized, alleviating some privacy concerns. Beyond these methodological advantages, social media itself has become an emotional outlet for many individuals, especially in East Asian contexts Yang and Li (2009); Zhou et al. (2023), where personal emotions are more freely expressed online rather than in face to face settings. Social media serves as a critical outlet for emotional expression, potentially offering richer and more authentic linguistic data for depression analysis.
However, most existing datasets are in English and focus solely on classification with binary or multi-class labels, lacking structured psychological insight or professional validation. Cai et al. (2023); Wang et al. (2020); Li et al. (2020). Moreover, with the emergence of generative models applied to mental health Hu et al. (2024); Xu et al. (2024); Lai et al. (2023); Yang et al. (2024), many scholars in psychology and computational linguistics now emphasize the need for datasets combining risk labels with structured analyses, such as user profiles or survey-based explanations. Such datasets align more closely with real-world settings, where multi-faceted insights are helpful for clinical intervention.
Some recent efforts have introduced summarization style datasets for depression, leveraging information extraction and automated summarization techniques Sotudeh et al. (2022, 2021). However, these datasets usually rely on model generated content validated by automatic metrics (e.g., ROUGE Lin (2004)) and typically lack professional mental health annotation, raising concerns about domain-specific reliability and accuracy.
To address these challenges, we introduce (CNSocialDepress, the first publicly available Chinese-language dataset for depression risk detection that integrates binary risk labels with structured psychological analyses. All annotations are drafted and validated by certified mental health experts, ensuring both domain relevance and annotation quality. CNSD is designed to support multiple task paradigms, including binary classification, generation of structured analyses, summarization, and fine-tuning of large language models (LLMs) for psychological reasoning.
Our contributions are fourfold:
  • We construct a high-quality Chinese dataset that bridges the gap between classification and interpretable analysis in depression detection, and supports a wide range of tasks including binary classification, generation of structured psychological analyses, summarization, and fine-tuning of large language models.
  • We design a robust annotation and generation pipeline involving professional experts and structured templates;
  • We conduct extensive experiments demonstrating that CNSD enables effective training and evaluation of LLM-based methods across diverse depression-related tasks.
  • We propose and validate a robust pipeline for generating structured psychological analyses in the context of depression risk (Section 4).
2 Related Work
2.1 Existing Datasets for Depression Detection
Current depression detection datasets primarily come from English-language social media platforms like Twitter, Reddit, Facebook, and Instagram Shen et al. (2017); Parapar et al. (2022); Zhang et al. (2021); Raihan et al. (2024). Annotation methods include self-disclosure identification Yates et al. (2017); Bathina et al. (2021); Islam et al. (2018), manual coding by clinical raters Almouzini et al. (2019); Yazdavar et al. (2020); Alhamed et al. (2024), and symptom mapping Seabrook et al. (2018); Aldarwish and Ahmad (2017); Zhang et al. (2022) based on DSM American Psychiatric Association et al. (2000) and PHQ Kroenke et al. (2001). While most datasets use binary classification, others incorporate severity scales or symptom-specific tagging Zhang et al. (2022); Mowery et al. (2017). Multilingual resources for depression detection exist for Spanish Romero et al. (2024), Arabic Maghraby and Ali (2022), Russian Stankevich et al. (2020), Portuguese Santos et al. (2024), Japanese Yuka Niimi (2021), and Thai Hämäläinen et al. (2021). For Chinese, Sina Weibo is the primary source Li et al. (2020); Shen et al. (2018); Li et al. (2023); Guo et al. (2023); Yang et al. (2021), with datasets like WU3D Wang et al. (2020) and SWDD Cai et al. (2023) commonly used.
2.2 Methodological Advancements
Early methods in depression detection used statistical approaches like TF-IDF and feature engineering Yang et al. (2020) along with traditional machine learning classifiers Cortes (1995); Li (2024); Breiman (2001); McCallum et al. (1998); Dreiseitl and Ohno-Machado (2002); Li et al. (2024); Shi et al. (2024). With the rise of neural networks Schmidhuber (2015), deep learning models such as Word2Vec Mikolov et al. (2013), LSTM Sak et al. (2014), CNN Kim (2014), Transformer Vaswani (2017), and BERT Devlin et al. (2019) improved the detection of depression. Recently, large language models (LLMs) have been applied to depression detection tasks due to their ability to generate labels and provide interpretable analyses Lan et al. (2024); Hu et al. (2024); Xin Yan (2023); Lai et al. (2023). These models enhance explainability in mental health analysis Wang et al. (2024c); Yang et al. (2024); Xu et al. (2024); Yang et al. (2023); Hu et al. (2024); Xu et al. (2024).
2.3 Summary and Gaps
Despite existing efforts, current datasets rarely include structured psychological analysis, especially for non-English social media content. Furthermore, available resources often lack professional mental-health annotation, limiting their practical utility. This study addresses these limitations by presenting the first expert-validated Chinese-language dataset featuring detailed structured annotations, thereby providing richer resources for depression-risk detection and analysis.
3 Construction of Dataset
Figure 1: Dataset Construction Process: During the data annotation process, we used a subset of the original SWDD dataset for labeling, which included 116 Positive (Depressive) users and 117 Negative users. Psychologists used rating scales based on the DSM-5 and the statistical results of the dataset’s text to formulate an initial labeling guideline. Then, during the psychologists’ labeling process, random sampling was continuously performed for cross validation, and the labeling standards were continuously updated based on the annotation outcomes. Ultimately, the labeled gold-standard data was obtained, with each user containing a six-dimensional structural analyses summary.
Our raw data are derived from the SWDD dataset Cai et al. (2023), a user-level dataset collected from Sina Weibo, the world’s largest Chinese-language social media platform. This dataset comprises two categories of users (depressed and non-depressed), each user has between dozens and hundreds of Weibo posts. The dataset, which is labeled as either depressed or non-depressed, also contains user-level depression feature annotations provided by experts in accordance with the DSM-5333https://www.psychiatry.org/psychiatrists/practice/dsm/educational-resources/dsm-5-fact-sheets American Psychiatric Association et al. (2000) reference standards.
From this SWDD dataset, we have 116 self-reported depressed (positive) users and 117 non-depressed (negative) users to form the candidate samples for our dataset. Each selected user has at least 60 posts, and the total token count for each user’s posts is no fewer than 3,000 tokens [^4] Qwen2.5’s tokenizer.
To guide the annotation process, a group of psychology experts formulated a preliminary annotation standard based on the DSM-5 diagnostic criteria, the PHQ-9 scale Kroenke et al. (2001), and prior research on linguistic features indicative of depression in online texts Mothe et al. (2022).
Criteria involving objective factual statements (e.g., clinical symptoms, medical records, self-reports) with stronger diagnostic specificity are designated as Primary Criteria, while criteria based on subjective linguistic or emotional expressions (e.g., negative phrasing, emotional intensity) with weaker specificity are designated as lower-weight Secondary Criteria, and the hierarchical classification follows clinical diagnostic principles that prioritize objective symptoms over subjective perceptions. This standard comprises six depression dimensions:
  • Dimension 1: Depressive Psychological State (primary criterion). At its core, this category reflects a loss of self-worth, overwhelming guilt, and suicidal ideation. Representative semantic segments include Inferiority, Sorry, Death, Self-harm, Wrist cutting, Suicide, etc.
  • Dimension 2: Depression-Related Medical Expressions (primary criterion). Focusing on expressions linked to medication usage and psychological diagnoses, this category covers terms such as Take medication, Venlafaxine, Side effects, Depression, Anxiety disorder, Hospital, Doctor, Disease, etc.
  • Dimension 3: Depression-Related Clinical Symptoms (primary criterion). Addressing physiological and somatic alterations, including changes in weight or appetite, disturbed sleep patterns, fatigue, and various forms of physical pain, this dimension includes semantic segments such as Appetite, Insomnia, Sleeping pills, Nightmare, Tired, Headache, Stomachache, among others.
  • Dimension 4: Negative Emotions (secondary criterion). Encompassing a broad array of adverse emotional states, this part presents examples such as Torment, Abyss, Hell, Sad, Grieve, Anxious, Depressed, Lonely, Despair, etc.
  • Dimension 5: Potential External Causes of Depression (secondary criterion). Highlighting possible triggers like marital or intimate relationship issues, family conflicts, significant social events, and everyday stressors, this segment comprises semantic units including Divorce, Parents, College entrance examination, School, Teacher, Poor, Drop out, etc.
  • Dimension 6: Depression-Related Patterns of Language Use (secondary criterion). Emphasizing particular linguistic patterns, such as negations, interrogatives, and derogatory expressions. This category includes illustrative semantic segments like I don’t know, I don’t like me, I really don’t want to, I can’t do it, why, and what should I do?, among others.
Figure 2: Example entry from the CNSD-Gold dataset
The annotation work was carried out by four senior researchers, each with substantial research experience in psychology, specializing in depression scales. Using an internally developed annotation platform, they identified specific depression dimensions present in each post, along with the corresponding semantic segments. Throughout this process, the annotators periodically performed random cross-validation to ensure reliability and consistency. Because this annotation task relies on qualitative rather than quantitative metrics, the annotation schema was updated as needed, informed by real-time feedback from the annotated texts and discussions among experts. During the annotation process, the expert annotators did not refer to the binary labels from the original dataset, consequently, some of the depression risk labels may differ from those in the original data set.
This rigorous process produced the CNSD Gold dataset (100 positive and 100 negative users) and the CNSD Test dataset (16 positive and 17 negative users). The dataset construction process can be referenced in Figure 1.
Each entry as shown in Figure 2 comprises two parts: a binary depression label and a detailed, dimension-specific analysis. For each dimension, the analysis includes an overall assessment and a list of post numbers with brief justifications for their selection.
The overall statistics are presented in Table 1, with a total of 10,306 dimension-level annotations summarized in Table 2.
Positive (Depressive) Negative (Non-depressive)
Number of Users 116 117
Number of Texts 20,360 23,818
Number of Tokens 1,024,978 1,115,951
Table 1: Dataset Statistics.
Dimension Negative User Positive User
Dim1 117 2127
Dim2 2 555
Dim3 126 768
Dim4 514 2933
Dim5 193 863
Dim6 231 1877
Total 1183 9123
Table 2: Dimension Annotation Statistics for Negative and Positive Users: This table shows the statistics of depression repeated segments marked by psychological experts across different dimensions. A text may contain multiple depression repeated segments.
4 Automated Dataset Generation Pipeline
Due to the substantial human effort and professional psychological expertise required for dataset annotation, user-level depression risk datasets based on social media remain scarce in the Chinese domain. To address this gap, we developed an automated dataset generation pipeline. This pipeline was used to create the CNSD Silver dataset by processing the SWDD binary classification dataset, leveraging insights from the manually annotated CNSD Gold dataset (described in Section 3). For the construction of CNSD Silver, we randomly selected 100 positive (depressive risk) users and 100 negative (non-depressive risk) users from the SWDD dataset. Each user contributed between several dozen and several hundred posts, totaling more than 3,000 tokens per user.
The pipeline consists of two main modules:
4.1 Module I: Dimension-Wise Automatic Labeling
This module uses a smaller model (Qwen2.5-14B Bai et al. (2023)) to automatically annotate texts across six depression-related dimensions ($\mathcal{D}=\{D_{1},D_{2},\dots,D_{6}\}$). The model is first fine-tuned on carefully curated text-level data derived from the positive users in CNSD Gold, enabling it to assign each individual text to one or more of these six dimensions.
Data Preparation for Fine-tuning Module I Model:
  1. 1.
    Texts from positive users in CNSD Gold containing depression-related semantic segments for each dimension $D_{k}$ were extracted, forming sets $S_{k}$.
  2. 2.
    The dimension with the fewest samples ($n_{\text{min}}=442$, corresponding to "depression-related medical expressions") was identified:
    $n_{\text{min}}=\min_{k\in\{1,2,\dots,6\}}n_{k}=442$
  3. 3.
    To balance data across dimensions, texts for each dimension $D_{k}$ were downsampled to $n_{\text{min}}$ if $n_{k}>n_{\text{min}}$, forming balanced sets $S^{\prime}_{k}$:
    $S^{\prime}_{k}=\begin{cases}S_{k},&\text{if }n_{k}=n_{\text{min}},\\ \text{RandomSubset}(S_{k},n_{\text{min}}),&\text{if }n_{k}>n_{\text{min}}.\end{cases}$
    This resulted in $6\times 442$ labeled positive texts. Each text is annotated in the format: “This text segment belongs to [category] because it references [feature],” supplemented by 20 randomly generated expressions sharing the same meaning (see Appendix A.2).
  4. 4.
    For negative data, posts from negative users in CNSD Gold were split, and texts with any depression-related dimension were filtered out. From the remaining unlabeled texts, $6\times 442$ were uniformly sampled and assigned one of 20 distinct label expressions indicating absence of negative emotions (details in Appendix A.3).
  5. 5.
    The Module I model (Qwen2.5-14B) was then fine-tuned using LoRA Hu et al. (2022) on this combined set of $2\times(6\times 442)$ texts for 1 epoch at a learning rate of 5e-05 (details in Appendix A.4).
4.2 Module II: Automatic Verification and Summarization
In the second module, we applied the fine-tuned model from Module I to label every text for each user, extracting all texts that contain any depression-related semantics. We then leveraged the Deepseek R1 671B model DeepSeek-AI et al. (2025) to verify the correctness of both labels and texts on a case-by-case basis. Following this verification step, we employed a few-shot approach to summarize the extracted texts. In the few-shot setting, we instruct the model to assume the role of experts in psychology and text sentiment analysis. Additionally, we provide the model with detailed task requirements and exemplary gold-standard answer samples. Details are in the Appendix 3.
5 Experiments
We design our experiments around two central research questions. First , how effectively our data generation pipeline in producing high-quality structured annotations? Second , how well do various large language models (LLMs) perform on tasks involving user-level depression risk summarization and structured psychological analysis?
Based on these questions, our experiments aim to:
  1. 1.
    To assess the quality of text generated for six-dimensional depression analysis by models fine-tuned on CNSD Gold and CNSD Silver, comparing them against baseline models and few-shot approaches.
  2. 2.
    To evaluate the performance of these fine-tuned models on the downstream task of user-level depression risk classification.
  3. 3.
    To benchmark the performance of various leading LLMs on a structured depression analysis summarization task using our CNSD Gold dataset.
Additionally, as shown in the Appendix 5.5, we demonstrate the use of the CNSD Gold dataset as the test set for the binary depression classification task, further illustrating the dataset’s versatility.
5.1 Baseline Models
In the following experiments, we employ DeepSeek-R1 Distill-14B DeepSeek-AI et al. (2025), Qwen2.5-14B QwenTeam (2024), GPT-4o  [^5] https://openai.com/index/hello-GPT-4o/, GPT-4o-Mini [^6] https://openai.com/index/GPT-4o-mini-advancing-cost-efficient-intelligence/, DeepSeek-R1-671B [^7] https://www.deepseek.com/ DeepSeek-AI et al. (2025), and Llama3-8B-Chinese-Chat Wang et al. (2024b) as our generative baseline models. We summarize the key characteristics of each model on A.1.
5.2 Experimental Setup
For our experiments, we used NVIDIA A800 80G and A100 80G GPUs. Unless otherwise specified, all generation experiments were conducted with a temperature setting of 0.7. All fine-tuning experiments were conducted using the LLama Factory framework Zheng et al. (2024), employing the LoRA method Hu et al. (2022) for one epoch with a learning rate of 5e-5.
Our experiments consist of two main parts. The first part evaluates the performance of our pipeline on data generation, while the second part presents benchmarks on CNSD Gold dataset for Structured Summarization and Analysis Generation.
In our classification tasks, we use accuracy, recall, precision, and F1-score as evaluation metrics. For text generation experiments, we employ BLEU Papineni et al. (2002), ROUGE-1 Lin (2004), and BERTScoreZhang et al. (2020) for performance assessment.
5.3 Task I: Data Generation
In this section, we conducted multiple experiments on the CNSD Test dataset. We primarily evaluated the Qwen2.5 14B models fine-tuned on CNSD Gold and CNSD Silver:
  • Qwen2.5 14B Gold: Fine-tuned using the CNSD Gold.
  • Qwen2.5 14B Silver: Fine-tuned using the CNSD Silver. which was generated by applying the pipeline described in Section 4 to 100 randomly selected positive users and 100 randomly selected negative users from the SWDD binary classification dataset.
We compared the performance of these models with other baselines in two aspects:
  1. 1.
    The performance of the depression risk classification.
  2. 2.
    The quality of the generated text in the six-dimensional analyses section (including text hallucinations).
Throughout these experiments, unless noted otherwise, we employed a user-level generation strategy. This approach involves providing the complete set of a user’s Weibo posts along with an instruction that requires the model to directly output the user’s depression label along with the corresponding justification.
Strategy BERTScore ROUGE-1 BLEU
Pipeline 0.791 0.478 0.288
FS:Qwen2.5 14B 0.649 0.076 0.075
FS:GPT-4o 0.678 0.269 0.094
FS:GPT-4o Mini 0.674 0.170 0.070
FS:DeepSeek R1 671B 0.678 0.301 0.054
FS:DeepSeek R1 Distill -14B 0.6756 0.191 0.073
Table 3: Comparison of Generation Strategies on Text Quality: In the table, FS stands for Few-Shot, and Pipeline is based on the Automated Dataset Generation Pipeline we proposed in Section 4 and our proposed dataset in Section 3. The prompt used in this experiment can be found in Appendix A.6. The purpose is to demonstrate that, using the same prompt, our pipeline achieves the highest text generation quality that meets our task requirements.
5.3.1 Six-Dimensional Analysis of Depression task
Research has demonstrated that structured knowledge enhances the benefits of integrating structured data into LLMs Moiseev et al. (2022), thereby improving output quality. CNSD Gold and CNSD Silver are datasets with fine-grained structured annotations across six dimensions to summarize and analyze social media users. Fine-tuning on these datasets should reduces textual hallucinations and achieves text generation quality comparable to the other much larger model.
Table 3 shows the comparisons between different generation strategies in terms of the quality of the produced text. Notably, our Pipeline achieves the highest scores across all three metrics:BERTScore (0.791), ROUGE-1 (0.478), and BLEU (0.288), indicating superior semantic alignment, lexical overlap, and n-gram precision, respectively. In contrast, with the FS (few-shot) strategies, all models: Qwen2.5 14B, GPT-4o, and DeepSeek R1 671B, yield comparatively lower performance on each metric. Even the large-scale DeepSeek R1 671B model does not surpass the Pipeline approach, suggesting that the proposed Pipeline method effectively leverages its generation procedure to produce higher-quality text that more closely matches the reference, outperforming both smaller and larger models under the few-shot setting.
Additionally, we also report the comparison of generation quality between the original model, Qwen2.5 14B Gold, and Qwen2.5 14B Silver. (Table 5)
For automatic metrics, the original model achieved a Bert-Score of 0.649, which increased to 0.764 for Gold (+16.9%) and 0.787 for Silver (+21.5%), showing stronger semantic alignment. Rouge-1 rose from 0.076 to 0.217 for Gold (+172%) and 0.218 for Silver (+172%), indicating improved lexical overlap. Bleu increased from 0.075 to 0.186 for Gold (+149%) and 0.237 for Silver (+218%), reflecting better n-gram precision and structural quality.
The human evaluation was conducted by two linguistic experts across accuracy, coverage, and hallucination. Compared to the original Qwen2.5 14B, both Silver and Gold showed clear improvements. Silver improved by about 33.4% in accuracy, 37.1% in coverage, and 15.5% in hallucination. Gold achieved even stronger gains, with 42.9% in accuracy, 58.0% in coverage, and 26.8% in hallucination. Overall, Gold performed best in human evaluation, while Silver, though slightly behind, still showed significant gains over the original and reached levels close to Gold.
In summary, Gold achieved the strongest results in human evaluation, while Silver performed better in automatic metrics. Both models clearly outperformed the original, showing that fine-tuning with our dataset or pipeline can significantly enhance generation quality at semantic, lexical, and structural levels.
Model/Test Dataset Swdd Origin Wu3d Ours (Gold)
Acc F1 Acc F1 Acc F1
RoBERTa Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.88 0.87
RoBERTa Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.82 0.78
Bert based Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.85 0.82
Bert based Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.83 0.80
StructBERT-mental (trained by swdd origin) 0.81 0.84 0.69 0.76 0.85 0.86
StructBERT-mental (trained by wu3d) 0.85 0.85 0.92 0.92 0.83 0.80
Llama3-8b 0.86 0.86 0.79 0.79 0.60 0.63
Glm4-9b-chat 0.76 0.81 0.77 0.81 0.79 0.82
Qwen2.5 7b 0.88 0.89 0.84 0.85 0.94 0.94
GPT-4o-Mini 0.65 0.74 0.65 0.74 0.53 0.68
GPT-4o 0.92 0.92 0.91 0.92 0.83 0.86
DeepSeek-R1 671b 0.86 0.87 0.89 0.90 0.81 0.84
Table 4: Model Classification Performance Comparison on Different Test Datasets.
Model Bert-Score Rouge1 Bleu H.acc H.cov H.hallu
Qwen2.5 14B 0.649 0.076 0.075 0.583 0.574 0.657
Qwen2.5 14B Silver 0.787 0.218 0.237 0.778 0.787 0.759
Qwen2.5 14B Gold 0.764 0.217 0.186 0.833 0.907 0.833
Table 5: Comparison of text quality across different models. ’Qwen2.5 14B Gold’ denotes the model fine-tuned on our proposed dataset, while ’Qwen2.5 14B Silver’ denotes the model fine-tuned using data generated from the pipeline strategy described in Table 3, with further details in Section 4. The table reports three human evaluation metrics: acc (accuracy of the generated text), cov (content coverage), and hallu (hallucination). Higher values across all metrics indicate better generation performance.
5.3.2 Classification task
In the depression classification task, Qwen2.5 14B Silver achieves the best performance compared to other models, with an accuracy of 0.944 and an F1 score of 0.941, even surpassing Qwen2.5 14B Gold, which was fine-tuned on human-annotated data. Furthermore, it outperforms large-scale models such as GPT-4o (accuracy = 0.917, F1 score = 0.923) and DeepSeek R1 671B (accuracy = 0.861, F1 score = 0.872). This result provides evidence for the effectiveness of our generation pipeline and the high quality of the data produced through this pipeline. The details of the experimental results are in the Table 6.
Model Accuracy Recall Precision F1
Qwen2.5 14B 0.889 0.944 0.850 0.894
Qwen2.5 14B_FT_Silver 0.944 0.889 1.000 0.941
Qwen2.5 14B_FT_Gold 0.861 0.889 0.842 0.864
GPT-4o 0.917 1.000 0.857 0.923
GPT-4o Mini 0.667 1.000 0.600 0.750
DeepSeek R1 671B 0.861 0.944 0.801 0.872
DeepSeek R1 Distill-14B 0.889 1.000 0.818 0.900
Table 6: Comparison of Models for Generated Dataset Quality: Application to Classification Tasks.
5.4 Task II: Structured Summarization and Analysis Generation
The proposed dataset effectively supports research on generative models in the field of internet depression risk analyses, demonstrating particular value for user depressive mental state summarization tasks. Due to the absence of comparable open-source datasets, we conducted preliminary evaluations on mainstream generative models using our annotated dataset containing 233 user-level samples. The detailed results are demonstrated in the table 7. We used the basic prompt as follows:
Please carefully read the following multiple texts and then answer the question: Based on these texts, does this user exhibit a depressive mood? If the user exhibits depressive mood, answer ’Yes’; if the user does not exhibit depressive mood, answer ’No’. Then, provide your step-by-step analysis process.
Model Bert-Score Rouge1 Bleu
Llama3-8b $0.654$ $0.194$ $0.122$
Glm4-9b-chat $0.694$ $0.208$ $0.140$
Qwen2.5-7b $0.680$ $0.208$ $0.093$
DeepSeek-R1-671B $0.698$ 0.454 $0.125$
GPT-4o-mini $0.696$ $0.166$ $0.126$
GPT-4o 0.710 $0.346$ 0.152
Table 7: Performance Comparison on Depression Risk Analysis Summarization Generation Task.
The results reveal variations across models in the task of risk analysis summarization generation:
Bert-Score: Narrow range (0.65-0.71) with GPT-4o leading at 0.71, indicating superior semantic understanding. Rouge1: DeepSeek-R1-671B achieves remarkable 0.45 score, suggesting 10% improvement over second-best performer. Bleu: Maximum 15.17 from GPT-4o vs minimum 9.36 from Qwen2.5-7b, reflecting lexical alignment differences.
5.5 Task III: Classification Experimentations
In the depression risk binary classification task, we used three datasets, including the original SWDD dataset Cai et al. (2023) and the WU3D dataset Wang et al. (2020). Both of these datasets were randomly sampled with 200 negative users and 200 positive users for the experiment. Details provided in the table 4.
In the classification experiments, in addition to the generative models mentioned in the section A.1, we also used models based on the BERT architecture with the details on A.7.
6 Conclusion
We introduced CNSocialDepress, a new dataset for depression risk analysis using Chinese social media texts. Uniquely, it is the first dataset to include both user-level and text-level labels, combining binary depression indicators with fine-grained six-dimensional analysis annotated by mental health professionals. We also developed an automated data generation pipeline to ease manual annotation and support language model fine-tuning for classification and summarization tasks. We believe CNSocialDepress will prove valuable across a range of applications, advancing early detection and intervention strategies in mental health.
7 Limitations
While CNSocialDepress provides a valuable resource for depression risk analyses in Chinese social media, several limitations should be noted. First, the dataset is sourced exclusively from Weibo, potentially introducing platform-specific demographic biases (e.g., underrepresentation of rural or elderly populations).
Second, linguistic diversity, including dialectal variations and metaphorical expressions common in Chinese depressive discourse, remains insufficiently captured and may limit cross-regional generalizability. Third, although manual expert annotation ensures high quality, it constrains scalability, resulting in a relatively limited dataset size.
Additionally, the current structured analysis focus primarily on depressive symptoms, omitting comorbid mental health conditions (e.g., anxiety) and contextual social factors. Future work will expand data collection to multi-platform sources, integrate dynamic lexicon updates for emerging expressions, and adopt hybrid annotation frameworks (e.g., expert-guided crowdsourcing) to enhance coverage and efficiency. Privacy-preserving techniques and longitudinal data tracking will further strengthen ethical and practical utility.
8 Ethic Statement
This dataset is built upon the SWDD Cai et al. (2023) benchmark released in early 2023. The original data comes from user-level text on the Weibo platform and was annotated by a team of psychology experts according to the DSM-5 symptom standards.
During our work on the re-annotation process, we applied a second round of deep anonymization to meet privacy requirements. This involved systematically removing personal information, geographic locations, and potential identifiers from the original text, ensuring compliance with data protection protocols Benton et al. (2017) and GDPR (General Data Protection Regulation [^8] EUR-Lex GDPR Regulation). We also invited experienced researchers specializing in depression-related psychology to provide professional annotations based on their expertise.
This dataset is mainly used for building early-warning models for depression risk. Through multi-dimensional feature analysis, it detects potential risk signals and can be used to automatically generate structured psychological assessment reports, offering technical support for digital mental health services. It is important to note that this study is strictly limited to the realm of mental health assistance, and any results should only serve as references for professional diagnosis or as a digital tool for self-assessment.
At present, none of the algorithmic models can replace in-person psychiatric evaluations or deliver pathological diagnoses. If a system built using this dataset flags high-risk cases, or if individuals already show persistent mood disturbances or impaired social functioning, they should seek professional medical attention promptly. Mental health issues require clinical expertise, and any online self-assessment tool cannot substitute for in-person medical diagnosis.
Appendix A Appendix
A.1 Baseline Models
  • DeepSeek-R1 is a large-scale model specialized in logical reasoning,This model excels in complex reasoning tasks such as mathematics and programming while maintaining strong overall performance on general tasks. Chosen for its specialized excellence in complex logical reasoning while maintaining strong general task performance, potentially beneficial for in-depth structured psychological analysis and nuanced depression risk detection.
  • DeepSeek R1 Distill-14B is a lightweight reasoning model derived from Qwen2.5-14B through Knowledge Distillation. Chosen for its lightweight reasoning capabilities, aiming to retain the advantages of larger models in complex tasks while reducing computational cost, making it suitable for efficient depression risk analysis.
  • Qwen2.5-14B is part of Alibaba’s Qwen series of large language models. Chosen for its strong performance in long-form text generation and instruction-following, relevant for generating structured psychological analyses from Chinese social media posts.
  • GPT-4o is a multimodal large language model released by OpenAI. Chosen for its state-of-the-art multimodal and multilingual processing capabilities, including Chinese, and its improved overall performance and response speed, making it a powerful baseline for depression detection and psychological analysis.
  • GPT-4o-Mini is a lightweight AI model developed by OpenAI to provide an efficient and cost-effective AI solution. Chosen as a resource- and cost-efficient model with robust text understanding and generation, suitable for less computationally demanding depression risk identification and basic psychological analysis tasks.
  • Llama3-8B-Chinese-Chat, derived from Meta-Llama-3-8B, is an instruction-tuned model optimized for Chinese dialogue tasks. It has been fine-tuned to process both Chinese and English inputs. Chosen for its optimization for Chinese dialogue tasks and ability to handle extended conversations, making it well-suited for analyzing conversational social media posts for depression risk and generating coherent psychological insights in Chinese.
A.2 20 Randomly Generated Expression for the Positive Users
  1. 1.
    Carefully read the complete user Weibo text below and extract key information across six specified dimensions according to the content. Also pay attention to the user’s depressive state, ensuring that each dimension’s explanation is consistent with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence exists in the text, provide the text index (e.g., text_xx, 文本_xx) and a brief note (e.g., text_xx[sad]). - If no evidence is found for a dimension, state clearly that no relevant evidence was found.
    Note: Only output text index and explanation, not the actual text.
  2. 2.
    Read the entire user Weibo text below and, based on its content, extract core information under the following six dimensions. Also consider the user’s depressive status, ensuring that each dimension’s description matches this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if relevant information is detected, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[unhappy]). - For dimensions without evidence, explicitly state no relevant evidence found.
    Note: Output should include only text index and explanation, not the actual text.
  3. 3.
    Read through the complete user Weibo text below and extract key information across six dimensions based on its content. Pay close attention to the user’s depressive state, ensuring that the explanations are consistent with it.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is found in a dimension, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[sorrow]). - If no evidence is found, indicate explicitly.
    Note: Provide only indices and explanations, not the original content.
  4. 4.
    Carefully read the full Weibo text below and extract key evidence across six dimensions according to the text. At the same time, take into account the user’s depressive state, ensuring that explanations for each dimension align with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if there is corresponding evidence, output the text index (e.g., text_xx, 文本_xx) and a related note (e.g., text_xx[frustrated]). - If no evidence is found, state that no evidence exists.
    Note: Only provide text indices and notes, not the raw text.
  5. 5.
    Please read the full Weibo text below and extract key information from six dimensions. At the same time, pay attention to the user’s depressive state, ensuring that your explanations match this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence is detected, output the text index (e.g., text_xx, 文本_xx) and a short description (e.g., text_xx[sadness]). - If no evidence is found, specify that no evidence was found.
    Note: Only output text indices and explanations, not the original text.
  6. 6.
    Read the complete Weibo text below carefully and extract key information across six dimensions based on the text. Pay attention to the user’s depressive state and ensure that your descriptions are consistent with it.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is found, provide the text index (e.g., text_xx, 文本_xx) and a brief description (e.g., text_xx[low mood]). - If no evidence is found, clearly state that no evidence exists.
    Note: Only provide text indices and explanations, not the original text.
  7. 7.
    Please read through the complete user Weibo text below and extract key information across six dimensions based on its content. Pay attention to the user’s depressive state, ensuring that each explanation is aligned with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence exists, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[loss]). - If no evidence is found, clearly state no relevant evidence.
    Note: Only provide text indices and notes, not the original text.
  8. 8.
    Read the Weibo text below and extract key evidence across six dimensions according to the content, while considering the user’s depressive state to ensure explanations are consistent.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If a dimension contains relevant evidence, provide the text index (e.g., text_xx, 文本_xx) and a short note (e.g., text_xx[low mood]). - If no evidence is found, state explicitly that no evidence exists.
    Note: Output only text indices and explanations, not the original text.
  9. 9.
    Read the complete Weibo text carefully and extract key evidence across six dimensions based on the text. Pay attention to the user’s depressive state, ensuring that explanations match this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if relevant evidence is found, output the text index (e.g., text_xx, 文本_xx) and a brief explanation (e.g., text_xx[depressed]). - If no evidence is found, state explicitly that there is none.
    Note: Provide only text indices and notes, not the original text.
  10. 10.
    Read through the entire Weibo text and extract core information across six dimensions according to the content. Pay attention to the user’s depressive state and ensure that the descriptions are consistent with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if relevant evidence exists, output the text index (e.g., text_xx, 文本_xx) and a short explanation (e.g., text_xx[pessimistic]). - If no evidence is found, state clearly that no evidence exists.
    Note: Only output text indices and explanations, not the original content.
  11. 11.
    Please read the user Weibo text below completely and extract essential information across six dimensions. Ensure that the explanations for each dimension align with the depressive state given.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is detected, list the text index (e.g., text_xx, 文本_xx) and a short description (e.g., text_xx[melancholy]). - If no evidence is found, specify that no evidence was detected.
    Note: Output should only contain indices and notes, not the actual text.
  12. 12.
    Read carefully the full Weibo text provided and extract key evidence from six dimensions. Pay attention to the user’s depressive state, ensuring that outputs remain consistent with this state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If a dimension contains evidence, output the text index (e.g., text_xx, 文本_xx) and a concise note (e.g., text_xx[disheartened]). - If no evidence exists, specify so.
    Note: Output indices and explanations only, not original content.
  13. 13.
    Please go through the complete Weibo text below and identify critical information across six dimensions. Ensure the notes correspond with the user’s depressive condition.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - Where evidence is found, list the text index (e.g., text_xx, 文本_xx) with a short explanation (e.g., text_xx[downhearted]). - If no evidence is detected, indicate explicitly.
    Note: Provide only text indices and notes, excluding original text.
  14. 14.
    Read carefully the following Weibo text and extract evidence across six preset dimensions. Make sure the explanation for each aligns with the depressive state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If evidence exists for a dimension, output the text index (e.g., text_xx, 文本_xx) and a short remark (e.g., text_xx[low mood]). - If no evidence is found, specify clearly.
    Note: Output only indices and remarks, not the raw text.
  15. 15.
    Read through the following Weibo text and extract major evidence across six given dimensions. Ensure the output is consistent with the depressive state specified.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If evidence is found, provide the text index (e.g., text_xx, 文本_xx) with a brief note (e.g., text_xx[hopeless]). - If no evidence appears, state clearly.
    Note: Only provide indices and notes, not original text.
  16. 16.
    Please carefully review the Weibo text below and extract key information from six preset dimensions. All notes must align with the depressive condition of the user.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If there is evidence, output the corresponding text index (e.g., text_xx, 文本_xx) and a short remark (e.g., text_xx[emotional slump]). - If not found, specify clearly that none exists.
    Note: Only include indices and notes, no original content.
  17. 17.
    Please go through the complete Weibo text and extract the critical information across six given dimensions. Make sure your notes are consistent with the depressive state provided.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - For each dimension, if evidence is found, provide the text index (e.g., text_xx, 文本_xx) with a concise explanation (e.g., text_xx[feeling low]). - If no evidence exists, indicate so.
    Note: Only output indices and explanations, not actual text.
  18. 18.
    Carefully read the Weibo text provided and extract important evidence across six specified dimensions. Keep the explanations aligned with the depressive state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If a dimension contains evidence, state the text index (e.g., text_xx, 文本_xx) and a short remark (e.g., text_xx[upset]). - If not, state that no evidence is found.
    Note: Output only indices and remarks, not text.
  19. 19.
    Read the Weibo text completely and extract relevant evidence from six dimensions. Ensure each dimension’s note is consistent with the depressive state given.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If relevant evidence is detected, provide the text index (e.g., text_xx, 文本_xx) and short description (e.g., text_xx[distressed]). - If not found, state clearly no evidence.
    Note: Only indices and notes, exclude text itself.
  20. 20.
    Please read through the following full Weibo text and extract essential evidence across six preset dimensions. Ensure all notes are consistent with the depressive state.
    [Input] User depressive state: {label}, Full Weibo text: {text}.
    [Output requirements] - If evidence is found, provide the text index (e.g., text_xx, 文本_xx) and a short explanation (e.g., text_xx[worry]). - If no evidence is detected, state that explicitly.
    Note: Only indices and explanations, not original text.
A.3 20 Distinct Label Expressions for the Negative Users
  1. 1.
    "This text does not contain any expressions of negative emotions, nor does it involve any dimensions of depression."
  2. 2.
    "There are no descriptions related to negative emotions in the text, so it does not fit any depression-related dimensions."
  3. 3.
    "This text does not include expressions of negative emotions, and therefore it is not classified under any depression dimensions."
  4. 4.
    "The content does not reflect negative emotions and does not belong to any depression indicators."
  5. 5.
    "This passage contains no expressions of negative emotions and is not applicable to any depression dimensions."
  6. 6.
    "There are no descriptions of negative emotions in the text, so it is not classified under any depression-related dimensions."
  7. 7.
    "This content does not exhibit negative emotions and therefore does not involve any depression dimensions."
  8. 8.
    "The text does not contain any expressions related to negative emotions, nor does it meet the criteria for depression dimensions."
  9. 9.
    "There are no signs of negative emotions in this text, so it does not belong to any depression dimension."
  10. 10.
    "The text does not describe negative emotions and does not cover any depression-related dimensions."
  11. 11.
    "This passage does not contain any expressions related to negative emotions and is not within the scope of depression dimensions."
  12. 12.
    "The content does not include expressions of negative emotions, so it does not fall into any depression dimensions."
  13. 13.
    "There are no expressions of negative emotions in the text, nor does it meet the criteria for depression dimensions."
  14. 14.
    "This text does not exhibit any negative emotions and does not involve any depression dimensions."
  15. 15.
    "This content does not include descriptions of negative emotions and therefore is not classified under any dimensions of depression."
  16. 16.
    "The text does not show any negative emotions and is not applicable to depression-related dimensions."
  17. 17.
    "No expressions of negative emotions are seen in this passage, so it does not meet any depression dimension standards."
  18. 18.
    "There are no expressions of negative emotions in the text, so it does not belong to any depression dimension."
  19. 19.
    "This text does not display any expressions related to negative emotions and does not cover the dimensions of depression."
  20. 20.
    "The content lacks descriptions of negative emotions and is therefore not classified under any depression dimensions."
A.4 Instructions for the Fine-Tuning Process
  • Complete Instruction = Task Instruction + Supplementary Instruction
Task Instruction:
This task involves the following 6 dimensions of depression:
  • Potential External Causes of Depression (Secondary Judgment Criterion)
  • Depression-Related Clinical Symptoms (Primary Judgment Criterion)
  • Depression-Related Language Expression Patterns (Secondary Judgment Criterion)
  • Depression-Related Medical Expressions (Primary Judgment Criterion)
  • Depressive Psychological State (Primary Judgment Criterion)
  • Negative Emotions (Secondary Judgment Criterion)
Supplementary Instructions (randomly select one):
  1. 1.
    "Please read the following text segment and determine whether it contains any of the above 6 expressions of depressive emotion, and provide a brief explanation; if none is present, please indicate that the text does not belong to any depression dimension."
  2. 2.
    "Read the following text and analyze whether it demonstrates any one of the aforementioned 6 depressive emotion expression dimensions, and provide a brief explanation; if such expression is absent, please indicate that the text does not correspond to any dimension."
  3. 3.
    "Please carefully read the following text segment and confirm whether any one of the 6 depressive emotion expression dimensions mentioned above exists, and include a brief explanation; if not, please state that the text does not belong to any dimension."
  4. 4.
    "Please read the following content and determine whether it embodies any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please indicate that the text does not cover any depression dimension."
  5. 5.
    "Read the following text segment and determine whether it contains any one of the 6 depressive emotion expressions mentioned above, and provide a concise explanation; if not, please state that the text does not belong to any depression dimension."
  6. 6.
    "Please read the following text and determine whether any one of the above 6 depressive emotion expression dimensions appears, and include a brief explanation; if not, please indicate that the text does not meet any dimension."
  7. 7.
    "Read the following text and check whether any one of the above 6 depressive emotion expression dimensions is presented, and provide a brief explanation; if not, please state that the text does not belong to any dimension."
  8. 8.
    "Please review the following text segment and determine whether any one of the above 6 depressive emotion expression dimensions exists, and provide a brief explanation; if not, then indicate that the text does not belong to any dimension."
  9. 9.
    "Read the following text and confirm whether any one of the above 6 depressive emotion expression dimensions is embodied, and provide a brief explanation; if not, please state that the text does not involve any depression dimension."
  10. 10.
    "Please read the following content and determine whether it contains any one of the 6 depressive emotion expression dimensions mentioned earlier, and provide a brief explanation; if not, please indicate that the text does not belong to any dimension."
  11. 11.
    "Read the following text segment and verify whether it demonstrates any one of the above 6 depressive emotion expression dimensions, and include a brief explanation; if not, please state that the text does not meet any dimension."
  12. 12.
    "Please read the following text and check whether it exhibits any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please indicate that the text does not cover any depression dimension."
  13. 13.
    "Read the following text and confirm whether any one of the 6 depressive emotion expressions mentioned above exists, and include a brief explanation; if such a feature is absent, please state that the text does not belong to any dimension."
  14. 14.
    "Please carefully read the following text segment and determine whether it embodies any one of the above 6 depressive emotion expression dimensions, and provide a concise explanation; if not, please indicate that the text does not correspond to any dimension."
  15. 15.
    "Read the following text segment and confirm whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please state that the text does not involve any depression dimension."
  16. 16.
    "Please read the following text segment and determine whether any one of the 6 depressive emotion expression dimensions mentioned above appears in the text, and include a brief explanation; if not, please indicate that the text does not belong to any dimension."
  17. 17.
    "Read the following content and check whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a brief explanation; if not, please state that the text does not belong to any depression dimension."
  18. 18.
    "Please read the following content and determine whether any one of the above 6 depressive emotion expression dimensions exists, and provide a brief explanation; if not, please indicate that the text does not involve any dimension."
  19. 19.
    "Read the following text segment and confirm whether it demonstrates any one of the above 6 depressive emotion expression dimensions mentioned above, and provide a brief explanation; if not, please state that the text does not meet any dimension."
  20. 20.
    "Please read the following text and determine whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a concise explanation; if not, please indicate that the text does not belong to any depression dimension."
A.5 Module II Prompt
Figure 3: Module II Prompt.
A.6 Prompt for Table 3 On Section 5.3.1
Prompt is as follows: You are a professional in psychology and text sentiment analysis. Please determine whether the user is at risk of depression based on the following six dimensions:
  • Negative emotions (secondary criterion)
  • Depressive psychological state (primary criterion)
  • Depression-related clinical symptoms (primary criterion)
  • Potential external factors causing depression (secondary criterion)
  • Depression-related medical expressions (primary criterion)
  • Depression-related language expression patterns (secondary criterion)
For each dimension, please cite the corresponding text reference and provide a concise explanation.
[Output Data Example]
  • Answer: “Yes” or “No.”
  • Briefly explain the manifestations in each of the six dimensions, and cite the corresponding text reference.
  • In conclusion, “The user is at risk of depression” or “The user is not at risk of depression.”
A.7 Bert type models
BERT-base-chinese [^9] https://huggingface.co/google-BERT/BERT-base-chinese is built on the BERT-Base architecture Devlin et al. (2019) and is pre-trained on extensive Chinese corpora. It adopts masked language modeling (MLM) and next sentence prediction (NSP) to learn deep bidirectional representations of Chinese text at the character level.
Chinese-RoBERTa-wwm-ext Cui et al. (2019) leverages RoBERTa Liu et al. (2019) with a whole word masking (WWM) approach. Instead of masking individual characters, it masks entire words during MLM, leading to a richer understanding of context. The extended pre-training (ext) on large Chinese corpora further refines its effectiveness in various NLP tasks.
StructBERT-mental Wang et al. (2019) is based on StructBERT, which incorporates linguistic structures into its pre-training. By focusing on word and sentence structure, and adding specific pre-training for mental health texts, it improves performance in identifying and categorizing mental health-related content.

### v2

@@
 Abstract
-Depression is a pressing global public health issue, yet publicly available Chinese-language resources for risk detection remain scarce and are mostly limited to binary classification.
-To address this limitation, we release CNSocialDepress, a benchmark dataset for depression risk detection from Chinese social media posts.
-The dataset contains 44,178 texts from 233 users, within which psychological experts annotated 10,306 depression-related segments.
-CNSocialDepress provides binary risk labels together with structured multi-dimensional psychological attributes, enabling interpretable and fine-grained analysis of depressive signals.
-Experimental results demonstrate its utility across a wide range of NLP tasks, including structured psychological profiling and fine-tuning of large language models for depression detection.
-Comprehensive evaluations highlight the dataset’s effectiveness and practical value for depression risk identification and psychological analysis, thereby providing insights to mental health applications tailored for Chinese-speaking populations.
+Depression is a pressing global public health issue, yet publicly available Chinese-language resources for depression risk detection remain scarce and largely focus on binary classification.
+To address this limitation, we release CNSocialDepress, a benchmark dataset for depression risk detection on Chinese social media.
+The dataset contains 44,178 posts from 233 users; psychological experts annotated 10,306 depression-related segments.
+CNSocialDepress provides binary risk labels along with structured, multidimensional psychological attributes, enabling interpretable and fine-grained analyses of depressive signals.
+Experimental results demonstrate the dataset’s utility across a range of NLP tasks, including structured psychological profiling and fine-tuning large language models for depression detection.
+Comprehensive evaluations highlight the dataset’s effectiveness and practical value for depression risk identification and psychological analysis, thereby providing insights for mental health applications tailored to Chinese-speaking populations.
+Keywords: Depression Detection, Chinese Social Media, Benchmark Dataset, Mental Health
+[…]
 CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis
-Jinyuan XU1,*, Tian LAN2,*, Xintao YU3, Xue HE3,4, Hezhi ZHANG5, Ying WANG6, Pierre Magistry1, Mathieu Valette1, Lei LI7,8†thanks: * These authors contributed equally to this work.thanks: † Corresponding Author.
-1 Ertim Inalco, 2 Milkuya Studio, 3 Sorbonne Université, 4 IRD Lab 5 Faculty of Psychology, Peking University 6 Faculty of Psychology and Cognitive Science, Beijing Normal University 7 University of Washington, 8 VitaSight
-1 Introduction
-Depressive disorders are among the most common mental health conditions worldwide, characterized by persistent low mood or loss of interest in daily activities.
-According to a 2023 report [^1] https://www.who.int/news-room/fact-sheets/detail/depression by the World Health Organization (WHO), approximately 280 million people globally suffer from depression.
+Jinyuan XU1*, Tian LAN2*, Xintao YU3, Xue HE3,4, Hezhi ZHANG5, Ying WANG6 Pierre Magistry1, Mathieu Valette1, Lei LI7†
+1Ertim Inalco, 2Milkuya Studio, 3Sorbonne Université, 4IRD Lab,
+5Faculty of Psychology, Peking University,
+6Faculty of Psychology and Cognitive Science, Beijing Normal University,
+7Beijing Institute of Technology
+Abstract content
+11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author.
+1.  Introduction
+Depressive disorders are among the most common mental health conditions worldwide and are characterized by persistent low mood or loss of interest in daily activities.
+According to a 2023 report [^1] https://www.who.int/news-room/fact-sheets/detail/depression by the World Health Organization (WHO), approximately 280 million people worldwide live with depression.
 In China, a 2024 survey by the Chinese Center for Disease Control and Prevention Wang et al.
-(2024a) estimates that around 95 million individuals are affected by depression, and out of the 280,000 suicides reported annually, 40% are linked to depressive disorders.
+(2024a) estimates that around 95 million individuals are affected by depression.
+Of the approximately 280,000 suicides reported annually in China, 40% are linked to depressive disorders.
 Moreover, research Wang et al.
-(2024a) shows that depression strongly correlates with both suicidal and non-suicidal self-injurious deaths.
-Motivated by the urgent need for the detection of depression, researchers increasingly apply machine learning (ML) and natural language processing (NLP) methods for automated analysis of depression risk Squires et al.
+(2024a) shows that depression is strongly associated with both suicidal behavior and non-suicidal self-injury.
+Motivated by the urgent need to detect depression, researchers increasingly apply machine learning (ML) and natural language processing (NLP) methods to automatically assess depression risk Squires et al.
 (2023); Hasib et al.
 (2023); Aleem et al.
-(2022).
-While early efforts have shown promising results, they are heavily constrained by the limitations of existing datasets.
+(2022); Liu et al.
+(2024); Shi et al.
+(2024); Jia and Li (2024).
+While early efforts have shown promising results, they remain constrained by the limitations of existing datasets.
 Traditional depression detection studies often rely on clinical data Bittar et al.
 (2019); Fernandes et al.
-(2018), or on transcripts from medical or psychological interviews resources Shen et al.
+(2018) or on transcripts from medical or psychological interviews Shen et al.
 (2022); Li et al.
-(2022) that are are expensive to collect, limited in size and diversity, and may not reflect the informal, emotionally nuanced expressions typical of real-world online environments.
-To overcome these limitations, recent research has shifted toward leveraging user-generated content on social media platforms [^2] https://github.com/bucuram/depression-datasets-nlp Harrigian et al.
-(2021), which provides rich linguistic signals and greater availability.
-Such data are also more easily anonymized, alleviating some privacy concerns.
-Beyond these methodological advantages, social media itself has become an emotional outlet for many individuals, especially in East Asian contexts Yang and Li (2009); Zhou et al.
-(2023), where personal emotions are more freely expressed online rather than in face to face settings.
-Social media serves as a critical outlet for emotional expression, potentially offering richer and more authentic linguistic data for depression analysis.
-However, most existing datasets are in English and focus solely on classification with binary or multi-class labels, lacking structured psychological insight or professional validation.
-Cai et al.
-(2023); Wang et al.
+(2022b), which are expensive to collect, limited in size and diversity, and may not reflect the informal, emotionally nuanced expressions typical of real-world online environments.
+To overcome these limitations, recent research has shifted toward leveraging user-generated content on social media platforms Bucur et al.
+(2025); Harrigian et al.
 (2021); Li et al.
+(2025d); Jiang et al.
+(2024); Cai et al.
+(2025a), which provides rich linguistic signals and is more readily available.
+Such data are also more easily anonymized, thereby alleviating some privacy concerns.
+Beyond these methodological advantages, social media has become an important channel for emotional expression, especially in East Asian contexts Zhou et al.
+(2023); Yang and Li (2009); Zhou et al.
+(2024); Yang et al.
+(2025), where personal emotions may be expressed more freely online than in face-to-face settings.
+As a result, social media can provide richer and more authentic linguistic signals for depression analysis.
+However, most existing datasets focus solely on classification with binary or multi-class labels and lack structured psychological insights or professional validation.
+Moreover, generative models are increasingly used in mental health applications Hu et al.
+(2024); Xu et al.
+(2024); Yang et al.
+(2024b); Lai et al.
+(2023); He and Qu (2025a); Gu et al.
+(2025); Cai et al.
+(2025b); Xu et al.
 (2025).
-Moreover, with the emergence of generative models applied to mental health Hu et al.
-(2024); Xu et al.
-(2024); Lai et al.
-(2023); Yang et al.
-(2024), many scholars in psychology and computational linguistics now emphasize the need for datasets combining risk labels with structured analyses, such as user profiles or survey-based explanations.
-Such datasets align more closely with real-world settings, where multi-faceted insights are helpful for clinical intervention.
-Some recent efforts have introduced summarization style datasets for depression, leveraging information extraction and automated summarization techniques Sotudeh et al.
-(2022, 2021).
-However, these datasets usually rely on model generated content validated by automatic metrics (e.g., ROUGE Lin (2004)) and typically lack professional mental health annotation, raising concerns about domain-specific reliability and accuracy.
-To address these challenges, we introduce (CNSocialDepress, the first publicly available Chinese-language dataset for depression risk detection that integrates binary risk labels with structured psychological analyses.
-All annotations are drafted and validated by certified mental health experts, ensuring both domain relevance and annotation quality.
-CNSD is designed to support multiple task paradigms, including binary classification, generation of structured analyses, summarization, and fine-tuning of large language models (LLMs) for psychological reasoning.
+Researchers in psychology and computational linguistics emphasize the need for datasets that combine risk labels with structured analyses, such as user profiles or survey-based explanations.
+Such datasets better match real-world settings.
+Multifaceted insights can support clinical decision-making and downstream intervention.
+Some recent efforts have introduced summarization-style datasets for depression.
+They leverage information extraction and automated summarization techniques Sotudeh et al.
+(2022, 2021); He and Qu (2025b); Shi et al.
+.
+However, these datasets often rely on model-generated content.
+The content is typically validated using automatic metrics such as ROUGE Lin (2004).
+They also lack annotation by mental health professionals, which raises concerns about domain-specific reliability and accuracy.
+To address these challenges, we introduce CNSocialDepress (CNSD), a publicly available Chinese-language dataset for depression risk detection that can be applied to the development of early-stage depression detection tools.
+It pairs binary risk labels with structured psychological analyses.
+All annotations are drafted and validated by certified mental health professionals, ensuring domain relevance and annotation quality.
+CNSD supports multiple task paradigms.
+These include binary classification, structured analysis generation, summarization, and fine-tuning large language models (LLMs) for psychological reasoning.
 Our contributions are fourfold:
-  • We construct a high-quality Chinese dataset that bridges the gap between classification and interpretable analysis in depression detection, and supports a wide range of tasks including binary classification, generation of structured psychological analyses, summarization, and fine-tuning of large language models.
-  • We design a robust annotation and generation pipeline involving professional experts and structured templates;
-  • We conduct extensive experiments demonstrating that CNSD enables effective training and evaluation of LLM-based methods across diverse depression-related tasks.
-  • We propose and validate a robust pipeline for generating structured psychological analyses in the context of depression risk (Section 4).
-2 Related Work
-2.1 Existing Datasets for Depression Detection
-Current depression detection datasets primarily come from English-language social media platforms like Twitter, Reddit, Facebook, and Instagram Shen et al.
+  • We release CNSD, a high-quality Chinese dataset for depression risk detection that combines binary labels with structured, interpretable psychological analyses.
+  • We propose an expert-in-the-loop annotation protocol with structured templates and quality control.
+  • We benchmark CNSD across diverse task settings, including classification, structured analysis generation, summarization, and LLM fine-tuning for psychological reasoning.
+  • We present and validate a pipeline for generating structured psychological analyses for depression risk (Section 4).
+2.  Related Work
+2.1.  Existing Datasets for Depression Detection
+Current datasets for depression detection mainly come from English-language social media platforms such as Twitter, Reddit, Facebook, and Instagram Shen et al.
 (2017); Parapar et al.
@@
 (2021); Raihan et al.
+(2024); Li et al.
+(2025c); Jiang et al.
 (2025).
-Annotation methods include self-disclosure identification Yates et al.
+Common annotation strategies include identifying self-disclosure Yates et al.
 (2017); Bathina et al.
@@
 (2018); Aldarwish and Ahmad (2017); Zhang et al.
-(2022) based on DSM American Psychiatric Association et al.
-(2000) and PHQ Kroenke et al.
+(2022); Li et al.
+(2025a).
+Symptom definitions are often based on DSM-5 criteria Association (2013) and questionnaire instruments such as the PHQ-9 Kroenke et al.
 (2001).
-While most datasets use binary classification, others incorporate severity scales or symptom-specific tagging Zhang et al.
+Most datasets use binary labels.
+Others incorporate severity scales or symptom-specific tags Zhang et al.
 (2022); Mowery et al.
+(2017); Guan et al.
 (2025).
-Multilingual resources for depression detection exist for Spanish Romero et al.
+Multilingual resources also exist for depression detection.
+They cover languages such as Spanish Romero et al.
 (2024), Arabic Maghraby and Ali (2022), Russian Stankevich et al.
@@
 (2021).
-For Chinese, Sina Weibo is the primary source Li et al.
+For Chinese, Sina Weibo is the most common data source Li et al.
 (2020); Shen et al.
@@
 (2023); Yang et al.
-(2021), with datasets like WU3D Wang et al.
-(2020) and SWDD Cai et al.
-(2023) commonly used.
-2.2 Methodological Advancements
-Early methods in depression detection used statistical approaches like TF-IDF and feature engineering Yang et al.
-(2020) along with traditional machine learning classifiers Cortes (1995); Li (2024); Breiman (2001); McCallum et al.
-(1998); Dreiseitl and Ohno-Machado (2002); Li et al.
-(2024); Shi et al.
+(2021); Li et al.
+(2025b).
+Datasets such as WU3D Wang et al.
+(2020a) and SWDD Cai et al.
+(2023) are widely used.
+$2.$  Methodological Advancements
+Early work relied on statistical representations such as TF–IDF and hand-crafted features Yang et al.
+(2020); Li et al.
+(2022a).
+These features were typically paired with traditional machine learning classifiers Cortes (1995); Breiman (2001); McCallum et al.
+(1998); Dreiseitl and Ohno-Machado (2002); He et al.
 (2026).
-With the rise of neural networks Schmidhuber (2015), deep learning models such as Word2Vec Mikolov et al.
-(2013), LSTM Sak et al.
-(2014), CNN Kim (2014), Transformer Vaswani (2017), and BERT Devlin et al.
-(2019) improved the detection of depression.
-Recently, large language models (LLMs) have been applied to depression detection tasks due to their ability to generate labels and provide interpretable analyses Lan et al.
+With the rise of neural approaches Schmidhuber (2015), representation learning became central to depression detection.
+Embedding methods such as Word2Vec Mikolov et al.
+(2013) and neural architectures including LSTMs Sak et al.
+(2014), CNNs Kim (2014), Transformers Vaswani (2017), and BERT Devlin et al.
+(2019) improved performance.
+Recently, large language models (LLMs) have been explored for depression detection Lan et al.
 (2024); Hu et al.
 (2024); Xin Yan (2023); Lai et al.
+(2023); Li et al.
+(2026); He and Qu (2025a); Shi et al.
 (2025).
-These models enhance explainability in mental health analysis Wang et al.
+They can produce predictions and generate natural-language rationales or structured analyses, which supports interpretability.
+LLM-based approaches have also been used to improve explainability in mental health analysis Wang et al.
 (2024c); Yang et al.
-(2024); Xu et al.
+(2024b); Xu et al.
 (2024); Yang et al.
 (2023); Hu et al.
-(2024); Xu et al.
+(2024); Gu et al.
 (2025).
-2.3 Summary and Gaps
-Despite existing efforts, current datasets rarely include structured psychological analysis, especially for non-English social media content.
-Furthermore, available resources often lack professional mental-health annotation, limiting their practical utility.
-This study addresses these limitations by presenting the first expert-validated Chinese-language dataset featuring detailed structured annotations, thereby providing richer resources for depression-risk detection and analysis.
-3 Construction of Dataset
-Figure 1: Dataset Construction Process: During the data annotation process, we used a subset of the original SWDD dataset for labeling, which included 116 Positive (Depressive) users and 117 Negative users.
-Psychologists used rating scales based on the DSM-5 and the statistical results of the dataset’s text to formulate an initial labeling guideline.
-Then, during the psychologists’ labeling process, random sampling was continuously performed for cross validation, and the labeling standards were continuously updated based on the annotation outcomes.
-Ultimately, the labeled gold-standard data was obtained, with each user containing a six-dimensional structural analyses summary.
-Our raw data are derived from the SWDD dataset Cai et al.
-(2023), a user-level dataset collected from Sina Weibo, the world’s largest Chinese-language social media platform.
-This dataset comprises two categories of users (depressed and non-depressed), each user has between dozens and hundreds of Weibo posts.
-The dataset, which is labeled as either depressed or non-depressed, also contains user-level depression feature annotations provided by experts in accordance with the DSM-5333https://www.psychiatry.org/psychiatrists/practice/dsm/educational-resources/dsm-5-fact-sheets American Psychiatric Association et al.
-(2000) reference standards.
-From this SWDD dataset, we have 116 self-reported depressed (positive) users and 117 non-depressed (negative) users to form the candidate samples for our dataset.
-Each selected user has at least 60 posts, and the total token count for each user’s posts is no fewer than 3,000 tokens [^4] Qwen2.5’s tokenizer.
-To guide the annotation process, a group of psychology experts formulated a preliminary annotation standard based on the DSM-5 diagnostic criteria, the PHQ-9 scale Kroenke et al.
-(2001), and prior research on linguistic features indicative of depression in online texts Mothe et al.
+2.3.  Summary and Gaps
+Despite these efforts, few datasets include structured psychological analyses, especially for non-English social media.
+Furthermore, available resources often lack annotation by mental health professionals, which limits practical utility.
+This study addresses these limitations by presenting an expert-validated Chinese-language dataset with detailed structured annotations.
+The dataset provides richer resources for depression risk detection and analysis.
+3.  Construction of Dataset
+Figure 1: Dataset construction process.
+We sample 116 positive (depressed) users and 117 negative users from SWDD for expert annotation.
+Psychologists draft an initial guideline based on DSM-5 and corpus statistics, and iteratively refine it through random cross-validation.
+The final gold-standard data include a six-dimensional structured analysis summary for each user.
+Our raw data come from the SWDD dataset Cai et al.
+(2023), a user-level corpus collected from Sina Weibo, one of the largest Chinese-language social media platforms.
+The dataset includes two user groups: depressed and non-depressed.
+Each user has dozens to hundreds of Weibo posts.
+In addition to binary user labels, SWDD provides expert-annotated user-level depression features based on DSM-5 criteria [^2] DSM-5 Fact Sheets.
+Association (2013).
+From SWDD, we select 116 self-reported depressed (positive) users who disclosed a clinical diagnosis of depression in their posts, and 117 non-depressed (negative) users as candidates.
+Each selected user has at least 60 posts.
+The total length of a user’s posts is at least 3,000 tokens, computed using the Qwen2.5 tokenizer.
+[^3] Qwen2.5 tokenizer.
+To guide annotation, a team of psychology experts drafted an initial guideline based on DSM-5 criteria, the PHQ-9 Kroenke et al.
+(2001), and prior work on linguistic markers of depression in online text Mothe et al.
 (2022).
-Criteria involving objective factual statements (e.g., clinical symptoms, medical records, self-reports) with stronger diagnostic specificity are designated as Primary Criteria, while criteria based on subjective linguistic or emotional expressions (e.g., negative phrasing, emotional intensity) with weaker specificity are designated as lower-weight Secondary Criteria, and the hierarchical classification follows clinical diagnostic principles that prioritize objective symptoms over subjective perceptions.
-This standard comprises six depression dimensions:
-  • Dimension 1: Depressive Psychological State (primary criterion).
-At its core, this category reflects a loss of self-worth, overwhelming guilt, and suicidal ideation.
-Representative semantic segments include Inferiority, Sorry, Death, Self-harm, Wrist cutting, Suicide, etc.
-  • Dimension 2: Depression-Related Medical Expressions (primary criterion).
-Focusing on expressions linked to medication usage and psychological diagnoses, this category covers terms such as Take medication, Venlafaxine, Side effects, Depression, Anxiety disorder, Hospital, Doctor, Disease, etc.
-  • Dimension 3: Depression-Related Clinical Symptoms (primary criterion).
-Addressing physiological and somatic alterations, including changes in weight or appetite, disturbed sleep patterns, fatigue, and various forms of physical pain, this dimension includes semantic segments such as Appetite, Insomnia, Sleeping pills, Nightmare, Tired, Headache, Stomachache, among others.
-  • Dimension 4: Negative Emotions (secondary criterion).
-Encompassing a broad array of adverse emotional states, this part presents examples such as Torment, Abyss, Hell, Sad, Grieve, Anxious, Depressed, Lonely, Despair, etc.
-  • Dimension 5: Potential External Causes of Depression (secondary criterion).
-Highlighting possible triggers like marital or intimate relationship issues, family conflicts, significant social events, and everyday stressors, this segment comprises semantic units including Divorce, Parents, College entrance examination, School, Teacher, Poor, Drop out, etc.
-  • Dimension 6: Depression-Related Patterns of Language Use (secondary criterion).
-Emphasizing particular linguistic patterns, such as negations, interrogatives, and derogatory expressions.
-This category includes illustrative semantic segments like I don’t know, I don’t like me, I really don’t want to, I can’t do it, why, and what should I do?, among others.
-Figure 2: Example entry from the CNSD-Gold dataset
-The annotation work was carried out by four senior researchers, each with substantial research experience in psychology, specializing in depression scales.
-Using an internally developed annotation platform, they identified specific depression dimensions present in each post, along with the corresponding semantic segments.
-Throughout this process, the annotators periodically performed random cross-validation to ensure reliability and consistency.
-Because this annotation task relies on qualitative rather than quantitative metrics, the annotation schema was updated as needed, informed by real-time feedback from the annotated texts and discussions among experts.
-During the annotation process, the expert annotators did not refer to the binary labels from the original dataset, consequently, some of the depression risk labels may differ from those in the original data set.
-This rigorous process produced the CNSD Gold dataset (100 positive and 100 negative users) and the CNSD Test dataset (16 positive and 17 negative users).
-The dataset construction process can be referenced in Figure 1.
-Each entry as shown in Figure 2 comprises two parts: a binary depression label and a detailed, dimension-specific analysis.
-For each dimension, the analysis includes an overall assessment and a list of post numbers with brief justifications for their selection.
-The overall statistics are presented in Table 1, with a total of 10,306 dimension-level annotations summarized in Table 2.
+We define two levels of criteria.
+Primary criteria cover objective statements with higher diagnostic specificity, such as clinical symptoms, medical records, and explicit self-reports.
+Secondary criteria capture subjective linguistic or emotional expressions with lower specificity, such as negative phrasing and emotional intensity.
+This hierarchy follows clinical principles that prioritize objective symptoms over subjective perceptions.
+The guidelines cover six dimensions:
+  • Dimension 1: Depressive psychological state (primary).
+This dimension reflects loss of self-worth, overwhelming guilt, and suicidal ideation.
+Representative spans include inferiority, apology, death, self-harm, and suicide.
+  • Dimension 2: Medical expressions related to depression (primary).
+This dimension covers mentions of medication use and clinical diagnoses, such as taking medication, venlafaxine, side effects, depression, anxiety disorder, hospital, and doctor.
+  • Dimension 3: Clinical symptoms related to depression (primary).
+This dimension captures physiological or somatic changes, including appetite or weight changes, sleep disturbance, fatigue, and physical pain.
+Representative spans include appetite, insomnia, sleeping pills, nightmares, tiredness, headache, and stomachache.
+  • Dimension 4: Negative emotions (secondary). This dimension covers adverse emotional states such as sadness, grief, anxiety, loneliness, and despair.
+  • Dimension 5: Potential external causes of depression (secondary).
+This dimension captures possible triggers such as intimate relationship issues, family conflict, major social events, and everyday stressors.
+Representative spans include divorce, parents, school, teacher, and dropping out.
+  • Dimension 6: Language Use Patterns Related to Depression (secondary).
+This dimension captures linguistic patterns such as negation, questions, and derogatory expressions.
+Representative spans include I don’t know, I don’t like myself, I don’t want to, I can’t do it, why, and what should I do ?.
+Figure 2: A user-level annotated entry from CNSD-Gold (English translations).
+Each user has multiple posts.
+We treat each post as the primary unit for review.
+Annotation is performed on semantic spans within each post.
+Each span is assigned to one of the six dimensions.
+A post may contain multiple spans, and different spans may belong to different dimensions.
+Annotation was conducted by four senior researchers with expertise in psychology and depression scales.
+Annotators were blinded to the original binary labels in SWDD to ensure independent judgment.
+They periodically cross-checked overlapping samples to maintain consistency.
+Because the task relied more on qualitative judgment than on fixed quantitative criteria, the annotation guideline was not set in stone from the outset.
+It was revised in rounds as experts reviewed newly labeled data and resolved disagreements through discussion, and earlier annotations were corrected when the updated standards called for it.
+Given this evolving process, conventional inter-annotator agreement scores are not directly applicable.
+Since our annotation follows a different scheme from SWDD and all labeling was performed independently of the original labels, some user-level labels in our dataset differ from those in the original dataset.
+This process yields the CNSD Gold set (100 positive and 100 negative users) and the CNSD Test set (16 positive and 17 negative users). Figure 1 summarizes the construction process.
+Each entry (Figure 2) contains a binary label and a dimension-specific analysis. For each dimension, the analysis provides an overall assessment. It also lists post indices with brief justifications.
+The overall statistics are presented in Table 1, and the 10,306 dimension-level annotations are summarized in Table 2.
 Positive (Depressive) Negative (Non-depressive)
-Number of Users 116 117
-Number of Texts 20,360 23,818
-Number of Tokens 1,024,978 1,115,951
+NO. of Users 116 117
+NO. of Texts 20,360 23,818
+NO. of Tokens 1,024,978 1,115,951
 Table 1: Dataset statistics.
@@
 Total 1183 9123
-Table 2: Dimension Annotation Statistics for Negative and Positive Users: This table shows the statistics of depression repeated segments marked by psychological experts across different dimensions.
-A text may contain multiple depression repeated segments.
-4 Automated Dataset Generation Pipeline
-Due to the substantial human effort and professional psychological expertise required for dataset annotation, user-level depression risk datasets based on social media remain scarce in the Chinese domain.
-To address this gap, we developed an automated dataset generation pipeline.
-This pipeline was used to create the CNSD Silver dataset by processing the SWDD binary classification dataset, leveraging insights from the manually annotated CNSD Gold dataset (described in Section 3).
-For the construction of CNSD Silver, we randomly selected 100 positive (depressive risk) users and 100 negative (non-depressive risk) users from the SWDD dataset.
-Each user contributed between several dozen and several hundred posts, totaling more than 3,000 tokens per user.
-The pipeline consists of two main modules:
-4.1 Module I: Dimension-Wise Automatic Labeling
-This module uses a smaller model (Qwen2.5-14B Bai et al.
-(2023)) to automatically annotate texts across six depression-related dimensions ($\mathcal{D}=\{D_{1},D_{2},\dots,D_{6}\}$).
-The model is first fine-tuned on carefully curated text-level data derived from the positive users in CNSD Gold, enabling it to assign each individual text to one or more of these six dimensions.
-Data Preparation for Fine-tuning Module I Model:
+Table 2: Dimension-level annotation statistics for negative and positive users.
+This table reports the number of expert-annotated depression-related spans in each dimension.
+A single text may contain multiple spans.
+4.  Automated Dataset Generation Pipeline
+Expert annotation requires substantial human effort and professional psychological expertise.
+As a result, user-level depression-risk datasets for Chinese social media remain scarce.
+To narrow this gap, we develop an automated dataset generation pipeline.
+We use it to construct CNSD Silver from the SWDD binary classification dataset.
+The pipeline is informed by the manually annotated CNSD Gold dataset (Section 3).
+To build CNSD Silver, we randomly sample 100 positive (depression-risk) users and 100 negative (non-risk) users from SWDD.
+Each user contributes dozens to hundreds of posts.
+We keep users with at least 3,000 tokens in total.
+The pipeline consists of two modules.
+4.1.  Module I: Dimension-Wise Automatic Labeling
+Module I uses a mid-sized model (Qwen2.5-14B Bai et al.
+(2023)) to automatically assign posts to the six depression-related dimensions ($\mathcal{D}=\{D_{1},D_{2},\dots,D_{6}\}$).
+We fine-tune the model on text-level training data curated from positive users in CNSD Gold.
+After fine-tuning, the model can assign each post to one or more dimensions.
+Data preparation for fine-tuning Module I.
   1. 1.
-    Texts from positive users in CNSD Gold containing depression-related semantic segments for each dimension $D_{k}$ were extracted, forming sets $S_{k}$.
+    For each dimension $D_{k}$, we extract posts from positive users in CNSD Gold that contain at least one annotated span of $D_{k}$. These posts form a set $S_{k}$.
   2. 2.
-    The dimension with the fewest samples ($n_{\text{min}}=442$, corresponding to "depression-related medical expressions") was identified:
-    $n_{\text{min}}=\min_{k\in\{1,2,\dots,6\}}n_{k}=442$
+    We identify the dimension with the fewest training posts. The minimum count is $n_{\text{min}}=442$, which corresponds to depression-related medical expressions:
+    $n_{\text{min}}=\min_{k\in\{1,2,\dots,6\}}n_{k}=442.$
   3. 3.
-    To balance data across dimensions, texts for each dimension $D_{k}$ were downsampled to $n_{\text{min}}$ if $n_{k}>n_{\text{min}}$, forming balanced sets $S^{\prime}_{k}$:
-    $S^{\prime}_{k}=\begin{cases}S_{k},&\text{if }n_{k}=n_{\text{min}},\\ \text{RandomSubset}(S_{k},n_{\text{min}}),&\text{if }n_{k}>n_{\text{min}}.\end{cases}$
-    This resulted in $6\times 442$ labeled positive texts.
-Each text is annotated in the format: “This text segment belongs to [category] because it references [feature],” supplemented by 20 randomly generated expressions sharing the same meaning (see Appendix A.2).
+    To balance dimensions, we downsample each $S_{k}$ to $n_{\text{min}}$ when $n_{k}>n_{\text{min}}$. This yields balanced sets $S^{\prime}_{k}$:
+    $\small S^{\prime}_{k}=\begin{cases}S_{k},&\text{if }n_{k}=n_{\text{min}},\\ \text{RandomSubset}(S_{k},n_{\text{min}}),&\text{if }n_{k}>n_{\text{min}}.\end{cases}$
+    We obtain $6\times 442$ labeled positive posts.
+Each post is formatted using an instruction template such as: “This post belongs to [category] because it mentions [evidence].” We further augment the labels with 20 paraphrased expressions with the same meaning (Appendix A.2).
   4. 4.
-    For negative data, posts from negative users in CNSD Gold were split, and texts with any depression-related dimension were filtered out.
-From the remaining unlabeled texts, $6\times 442$ were uniformly sampled and assigned one of 20 distinct label expressions indicating absence of negative emotions (details in Appendix A.3).
+    For negative training data, we split posts from negative users in CNSD Gold into individual texts.
+We remove any text that contains depression-related content in any dimension.
+From the remaining texts, we uniformly sample $6\times 442$ examples.
+We assign them negative label expressions indicating that there is no evidence for the target dimension (As with the positive examples, we create 20 negative label expressions to reduce overfitting; details are provided in A.3).
   5. 5.
-    The Module I model (Qwen2.5-14B) was then fine-tuned using LoRA Hu et al.
-(2022) on this combined set of $2\times(6\times 442)$ texts for 1 epoch at a learning rate of 5e-05 (details in Appendix A.4).
-4.2 Module II: Automatic Verification and Summarization
-In the second module, we applied the fine-tuned model from Module I to label every text for each user, extracting all texts that contain any depression-related semantics.
-We then leveraged the Deepseek R1 671B model DeepSeek-AI et al.
-(2025) to verify the correctness of both labels and texts on a case-by-case basis.
-Following this verification step, we employed a few-shot approach to summarize the extracted texts.
-In the few-shot setting, we instruct the model to assume the role of experts in psychology and text sentiment analysis.
-Additionally, we provide the model with detailed task requirements and exemplary gold-standard answer samples.
-Details are in the Appendix 3.
-5 Experiments
-We design our experiments around two central research questions.
-First , how effectively our data generation pipeline in producing high-quality structured annotations?
-Second , how well do various large language models (LLMs) perform on tasks involving user-level depression risk summarization and structured psychological analysis?
-Based on these questions, our experiments aim to:
+    We fine-tune Qwen2.5-14B with LoRA Hu et al.
+(2022) on the combined set of $2\times(6\times 442)$ texts for 1 epoch.
+We use a learning rate of $5\times 10^{-5}$, with the prompt provided in Appendix A.4.
+4.2.  Module II: Automatic Verification and Summarization
+In Module II, we apply the fine-tuned Module I model to label all posts for each user.
+We then collect posts that are assigned to at least one dimension.
+Next, we use DeepSeek-R1-671B DeepSeek-AI et al.
+(2025) to verify the assigned labels and the supporting evidence on a case-by-case basis.
+After verification, we summarize the collected posts using a few-shot prompting setup.
+The prompt instructs the model to act as a psychology expert and to produce structured summaries aligned with CNSD Gold.
+We also provide detailed requirements and gold-standard examples.
+The full prompt is shown in Appendix A.7.
+5.  Experiments
+We design our experiments to answer two research questions.
+First, how effectively does our data generation pipeline produce high-quality structured annotations ?
+Second, how well do various large language models (LLMs) perform on user-level depression risk summarization and structured psychological analysis ?
+Based on these questions, we conduct three sets of experiments:
   1. 1.
-    To assess the quality of text generated for six-dimensional depression analysis by models fine-tuned on CNSD Gold and CNSD Silver, comparing them against baseline models and few-shot approaches.
+    To assess the quality of model-generated six-dimensional structured depression analyses. We compare models fine-tuned on CNSD Gold and CNSD Silver with baseline models and few-shot prompting.
   2. 2.
-    To evaluate the performance of these fine-tuned models on the downstream task of user-level depression risk classification.
+    To evaluate the performance of these fine-tuned models on user-level depression risk classification.
   3. 3.
-    To benchmark the performance of various leading LLMs on a structured depression analysis summarization task using our CNSD Gold dataset.
-Additionally, as shown in the Appendix 5.5, we demonstrate the use of the CNSD Gold dataset as the test set for the binary depression classification task, further illustrating the dataset’s versatility.
-5.1 Baseline Models
-In the following experiments, we employ DeepSeek-R1 Distill-14B DeepSeek-AI et al.
+    To benchmark leading LLMs on a structured depression analysis summarization task using CNSD Gold.
+In Section 5.5, we also use CNSD Gold as a test set for binary depression classification. This further illustrates the dataset’s versatility.
+5.1.  Baseline Models
+We use the following generative baselines: DeepSeek-R1-Distill-14B DeepSeek-AI et al.
 (2025), Qwen2.5-14B QwenTeam (2024), GPT-4o [^4] https://openai.com/index/hello-GPT-4o/, GPT-4o-mini [^5] https://openai.com/index/GPT-4o-mini-advancing-cost-efficient-intelligence/, DeepSeek-R1-671B [^6] https://www.deepseek.com/ DeepSeek-AI et al.
 (2025), and Llama3-8B-Chinese-Chat Wang et al.
-(2024b) as our generative baseline models.
-We summarize the key characteristics of each model on A.1.
-5.2 Experimental Setup
-For our experiments, we used NVIDIA A800 80G and A100 80G GPUs.
-Unless otherwise specified, all generation experiments were conducted with a temperature setting of 0.7.
-All fine-tuning experiments were conducted using the LLama Factory framework Zheng et al.
-(2024), employing the LoRA method Hu et al.
-(2022) for one epoch with a learning rate of 5e-5.
-Our experiments consist of two main parts.
-The first part evaluates the performance of our pipeline on data generation, while the second part presents benchmarks on CNSD Gold dataset for Structured Summarization and Analysis Generation.
-In our classification tasks, we use accuracy, recall, precision, and F1-score as evaluation metrics.
-For text generation experiments, we employ BLEU Papineni et al.
-(2002), ROUGE-1 Lin (2004), and BERTScoreZhang et al.
-(2020) for performance assessment.
-5.3 Task I: Data Generation
-In this section, we conducted multiple experiments on the CNSD Test dataset. We primarily evaluated the Qwen2.5 14B models fine-tuned on CNSD Gold and CNSD Silver:
-  • Qwen2.5 14B Gold: Fine-tuned using the CNSD Gold.
-  • Qwen2.5 14B Silver: Fine-tuned using the CNSD Silver.
-which was generated by applying the pipeline described in Section 4 to 100 randomly selected positive users and 100 randomly selected negative users from the SWDD binary classification dataset.
-We compared the performance of these models with other baselines in two aspects:
+(2024b).
+We summarize key characteristics in Table A.1.
+5.2.  Experimental Setup
+We use NVIDIA A800 (80GB) and A100 (80GB) GPUs.
+Unless otherwise specified, we set the generation temperature to 0.7.
+All fine-tuning experiments use the LLaMA-Factory framework Zheng et al.
+(2024).
+We apply LoRA Hu et al.
+(2022) for 1 epoch with a learning rate of $5\times 10^{-5}$.
+Our experiments have two parts. The first evaluates the pipeline for data generation. The second benchmarks models on the CNSD Gold dataset for structured summarization and analysis generation.
+For classification, we report accuracy, precision, recall, and F1-score. For generation, we report BLEU Papineni et al. (2002), ROUGE-1 Lin (2004), and BERTScore Zhang et al. (2020).
+5.3.  Task I: Data Generation
+We conduct experiments on the CNSD Test dataset. We focus on two Qwen2.5-14B models fine-tuned on CNSD Gold and CNSD Silver:
+  • Qwen2.5-14B Gold: Fine-tuned on CNSD Gold.
+  • Qwen2.5-14B Silver: Fine-tuned on CNSD Silver. CNSD Silver is generated using the pipeline in Section 4 on 100 positive and 100 negative users sampled from SWDD.
+We compare these models with other baselines in two aspects:
   1. 1.
-    The performance of the depression risk classification.
+    Depression risk classification performance.
   2. 2.
-    The quality of the generated text in the six-dimensional analyses section (including text hallucinations).
-Throughout these experiments, unless noted otherwise, we employed a user-level generation strategy.
-This approach involves providing the complete set of a user’s Weibo posts along with an instruction that requires the model to directly output the user’s depression label along with the corresponding justification.
+    Quality of the generated six-dimensional analyses, including hallucinated or unsupported content.
+Unless noted otherwise, we use user-level generation. We concatenate all posts from a user as input. We instruct the model to output a binary label and a corresponding justification.
 Strategy BERTScore ROUGE-1 BLEU
@@
 FS:DeepSeek R1 Distill -14B 0.6756 0.191 0.073
-Table 3: Comparison of Generation Strategies on Text Quality: In the table, FS stands for Few-Shot, and Pipeline is based on the Automated Dataset Generation Pipeline we proposed in Section 4 and our proposed dataset in Section 3.
-The prompt used in this experiment can be found in Appendix A.6.
-The purpose is to demonstrate that, using the same prompt, our pipeline achieves the highest text generation quality that meets our task requirements.
-5.3.1 Six-Dimensional Analysis of Depression task
-Research has demonstrated that structured knowledge enhances the benefits of integrating structured data into LLMs Moiseev et al.
-(2022), thereby improving output quality.
-CNSD Gold and CNSD Silver are datasets with fine-grained structured annotations across six dimensions to summarize and analyze social media users.
-Fine-tuning on these datasets should reduces textual hallucinations and achieves text generation quality comparable to the other much larger model.
-Table 3 shows the comparisons between different generation strategies in terms of the quality of the produced text.
-Notably, our Pipeline achieves the highest scores across all three metrics:BERTScore (0.791), ROUGE-1 (0.478), and BLEU (0.288), indicating superior semantic alignment, lexical overlap, and n-gram precision, respectively.
-In contrast, with the FS (few-shot) strategies, all models: Qwen2.5 14B, GPT-4o, and DeepSeek R1 671B, yield comparatively lower performance on each metric.
-Even the large-scale DeepSeek R1 671B model does not surpass the Pipeline approach, suggesting that the proposed Pipeline method effectively leverages its generation procedure to produce higher-quality text that more closely matches the reference, outperforming both smaller and larger models under the few-shot setting.
-Additionally, we also report the comparison of generation quality between the original model, Qwen2.5 14B Gold, and Qwen2.5 14B Silver. (Table 5)
-For automatic metrics, the original model achieved a Bert-Score of 0.649, which increased to 0.764 for Gold (+16.9%) and 0.787 for Silver (+21.5%), showing stronger semantic alignment.
-Rouge-1 rose from 0.076 to 0.217 for Gold (+172%) and 0.218 for Silver (+172%), indicating improved lexical overlap.
-Bleu increased from 0.075 to 0.186 for Gold (+149%) and 0.237 for Silver (+218%), reflecting better n-gram precision and structural quality.
-The human evaluation was conducted by two linguistic experts across accuracy, coverage, and hallucination.
-Compared to the original Qwen2.5 14B, both Silver and Gold showed clear improvements.
-Silver improved by about 33.4% in accuracy, 37.1% in coverage, and 15.5% in hallucination.
-Gold achieved even stronger gains, with 42.9% in accuracy, 58.0% in coverage, and 26.8% in hallucination.
-Overall, Gold performed best in human evaluation, while Silver, though slightly behind, still showed significant gains over the original and reached levels close to Gold.
-In summary, Gold achieved the strongest results in human evaluation, while Silver performed better in automatic metrics.
-Both models clearly outperformed the original, showing that fine-tuning with our dataset or pipeline can significantly enhance generation quality at semantic, lexical, and structural levels.
-Model/Test Dataset Swdd Origin Wu3d Ours (Gold)
-Acc F1 Acc F1 Acc F1
-RoBERTa Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.88 0.87
-RoBERTa Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.82 0.78
-Bert based Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.85 0.82
-Bert based Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.83 0.80
-StructBERT-mental (trained by swdd origin) 0.81 0.84 0.69 0.76 0.85 0.86
-StructBERT-mental (trained by wu3d) 0.85 0.85 0.92 0.92 0.83 0.80
-Llama3-8b 0.86 0.86 0.79 0.79 0.60 0.63
-Glm4-9b-chat 0.76 0.81 0.77 0.81 0.79 0.82
-Qwen2.5 7b 0.88 0.89 0.84 0.85 0.94 0.94
-GPT-4o-Mini 0.65 0.74 0.65 0.74 0.53 0.68
-GPT-4o 0.92 0.92 0.91 0.92 0.83 0.86
-DeepSeek-R1 671b 0.86 0.87 0.89 0.90 0.81 0.84
-Table 4: Model Classification Performance Comparison on Different Test Datasets.
-Model Bert-Score Rouge1 Bleu H.acc H.cov H.hallu
-Qwen2.5 14B 0.649 0.076 0.075 0.583 0.574 0.657
-Qwen2.5 14B Silver 0.787 0.218 0.237 0.778 0.787 0.759
-Qwen2.5 14B Gold 0.764 0.217 0.186 0.833 0.907 0.833
-Table 5: Comparison of text quality across different models.
-’Qwen2.5 14B Gold’ denotes the model fine-tuned on our proposed dataset, while ’Qwen2.5 14B Silver’ denotes the model fine-tuned using data generated from the pipeline strategy described in Table 3, with further details in Section 4.
-The table reports three human evaluation metrics: acc (accuracy of the generated text), cov (content coverage), and hallu (hallucination).
-Higher values across all metrics indicate better generation performance.
-5.3.2 Classification task
-In the depression classification task, Qwen2.5 14B Silver achieves the best performance compared to other models, with an accuracy of 0.944 and an F1 score of 0.941, even surpassing Qwen2.5 14B Gold, which was fine-tuned on human-annotated data.
-Furthermore, it outperforms large-scale models such as GPT-4o (accuracy = 0.917, F1 score = 0.923) and DeepSeek R1 671B (accuracy = 0.861, F1 score = 0.872).
-This result provides evidence for the effectiveness of our generation pipeline and the high quality of the data produced through this pipeline.
-The details of the experimental results are in the Table 6.
+Table 3: Comparison of generation strategies on text quality.
+In the table, FS stands for Few-Shot.
+Pipeline refers to the automated dataset generation pipeline proposed in Section 4, applied to our proposed dataset in Section 3.
+The purpose is to show that, with the same prompt, our pipeline achieves the highest text generation quality while meeting our task requirements.
+The prompt used in this experiment is shown in Figure 3.
+Figure 3: Prompt used for Table 3.
+5.3.1.  Six-Dimensional Structured Depression Analysis
+Prior work suggests that incorporating structured knowledge can improve LLM outputs Moiseev et al.
+(2022).
+CNSD Gold and CNSD Silver provide fine-grained structured annotations across six dimensions.
+They support user-level summarization and analysis of social media content.
+We hypothesize that fine-tuning on these datasets can reduce hallucinations.
+It may also improve generation quality compared to few-shot prompting and larger models.
+Table 3 compares different generation strategies using automatic metrics.
+Our Pipeline achieves the best scores on all three metrics: BERTScore (0.791), ROUGE-1 (0.478), and BLEU (0.288).
+These results indicate stronger semantic alignment, higher lexical overlap, and better $n$-gram precision.
+Few-shot (FS) prompting yields lower scores across models, including Qwen2.5-14B, GPT-4o, and DeepSeek-R1-671B.
+Notably, DeepSeek-R1-671B does not outperform the Pipeline.
+This suggests that the proposed Pipeline produces text that matches the reference more closely than few-shot prompting, across both smaller and larger models.
+We also compare the original Qwen2.5-14B model with Qwen2.5-14B Gold and Qwen2.5-14B Silver (Table 4). We use two complementary evaluation methods.
+  • Automatic metrics.
+The original model achieves a BERTScore of 0.649.
+It increases to 0.764 for Gold (+16.9%) and 0.787 for Silver (+21.5%).
+ROUGE-1 rises from 0.076 to 0.217 for Gold (+172%) and 0.218 for Silver (+172%).
+BLEU improves from 0.075 to 0.186 for Gold (+149%) and 0.237 for Silver (+218%).
+  • Human evaluation.
+Two linguistic experts evaluate accuracy, coverage, and hallucination.
+Compared to the original Qwen2.5-14B, both Gold and Silver improve substantially.
+Silver increases by 33.4% in accuracy and 37.1% in coverage.
+Its hallucination score increases by 15.5% (higher is better).
+Gold achieves larger gains.
+Accuracy increases by 42.9% and coverage increases by 58.0%.
+The hallucination score increases by 26.8%.
+Overall, Gold performs best in human evaluation.
+Silver is slightly behind but remains close to Gold and clearly outperforms the original model.
+In summary, Gold performs best in human evaluation.
+Silver performs better on automatic metrics.
+Both models substantially outperform the original model.
+These results show that fine-tuning with our dataset and pipeline can improve generation quality at semantic, lexical, and structural levels.
+Model BERTScore ROUGE-1 BLEU Human.acc Human.cov Human.hallu
+Qwen2.5-14B 0.649 0.076 0.075 0.583 0.574 0.657
+Qwen2.5-14B Silver 0.787 0.218 0.237 0.778 0.787 0.759
+Qwen2.5-14B Gold 0.764 0.217 0.186 0.833 0.907 0.833
+Table 4: Comparison of generation quality across models.
+Qwen2.5-14B Gold is fine-tuned on CNSD Gold.
+Qwen2.5-14B Silver is fine-tuned on CNSD Silver, which is generated by our pipeline (Section 4).
+We report automatic metrics (BERTScore, ROUGE-1, BLEU) and human evaluation metrics: acc (accuracy), cov (content coverage), and hallu (hallucination; higher is better).
+Higher values indicate better performance for all metrics.
+5.3.2.  Classification Task
+In the depression classification task, Qwen2.5-14B Silver achieves the best performance among all models.
+It reaches an accuracy of 0.944 and an F1 score of 0.941.
+It also surpasses Qwen2.5-14B Gold, which is fine-tuned on human-annotated data.
+Moreover, it outperforms large-scale models such as GPT-4o (accuracy = 0.917, F1 = 0.923) and DeepSeek-R1-671B (accuracy = 0.861, F1 = 0.872).
+These results support the effectiveness of our generation pipeline and suggest that the resulting silver data are of high quality.
+Detailed results are reported in Table 5.
 Model Accuracy Recall Precision F1
-Qwen2.5 14B 0.889 0.944 0.850 0.894
-Qwen2.5 14B_FT_Silver 0.944 0.889 1.000 0.941
-Qwen2.5 14B_FT_Gold 0.861 0.889 0.842 0.864
+Qwen2.5-14B 0.889 0.944 0.850 0.894
+Qwen2.5-14B (FT, Silver) 0.944 0.889 1.000 0.941
+Qwen2.5-14B (FT, Gold) 0.861 0.889 0.842 0.864
 GPT-4o 0.917 1.000 0.857 0.923
-GPT-4o Mini 0.667 1.000 0.600 0.750
-DeepSeek R1 671B 0.861 0.944 0.801 0.872
-DeepSeek R1 Distill-14B 0.889 1.000 0.818 0.900
+GPT-4o-mini 0.667 1.000 0.600 0.750
+DeepSeek-R1-671B 0.861 0.944 0.801 0.872
+DeepSeek-R1-Distill-14B 0.889 1.000 0.818 0.900
 Table 5: Comparison of Models for Generated Dataset Quality: Application to Classification Tasks.
-5.4 Task II: Structured Summarization and Analysis Generation
-The proposed dataset effectively supports research on generative models in the field of internet depression risk analyses, demonstrating particular value for user depressive mental state summarization tasks.
-Due to the absence of comparable open-source datasets, we conducted preliminary evaluations on mainstream generative models using our annotated dataset containing 233 user-level samples.
-The detailed results are demonstrated in the table 7.
-We used the basic prompt as follows:
-Please carefully read the following multiple texts and then answer the question: Based on these texts, does this user exhibit a depressive mood?
-If the user exhibits depressive mood, answer ’Yes’; if the user does not exhibit depressive mood, answer ’No’.
-Then, provide your step-by-step analysis process.
-Model Bert-Score Rouge1 Bleu
+5.4.  Task II: Structured Summarization and Analysis Generation
+Our dataset supports research on generative models for depression risk assessment from social media.
+It is particularly useful for user-level summarization of depressive signals.
+To the best of our knowledge, there is no directly comparable open-source dataset with the same annotation schema.
+We therefore conduct an initial benchmark of mainstream generative models on our annotated dataset of 233 users.
+Results are reported in Table 6.
+We use the following prompt:
+  • Instruction: Read the following texts written by one user. Then answer the question.
+  • Question: Based on these texts, does this user exhibit a depressive mood ?
+    • If the user exhibits a depressive mood, answer Yes.
+    • If the user does not exhibit a depressive mood, answer No.
+  • Output requirements: Provide a brief justification based on evidence from the texts. Then produce a structured summary aligned with the annotation schema.
+The results vary across models.
+BERTScore falls in a narrow range (0.654–0.710), with GPT-4o achieving the highest score.
+ROUGE-1 is highest for DeepSeek-R1-671B (0.454), which exceeds the second-best model (GPT-4o at 0.346) by an absolute margin of 0.108.
+BLEU ranges from 0.093 (Qwen2.5-7B) to 0.152 (GPT-4o).
+Model BERT Score ROUGE-1 BLEU
 Llama3-8B 0.654 0.194 0.122
@@
 GPT-4o 0.710 0.346 0.152
-Table 7: Performance Comparison on Depression Risk Analysis Summarization Generation Task.
-The results reveal variations across models in the task of risk analysis summarization generation:
-Bert-Score: Narrow range (0.65-0.71) with GPT-4o leading at 0.71, indicating superior semantic understanding.
-Rouge1: DeepSeek-R1-671B achieves remarkable 0.45 score, suggesting 10% improvement over second-best performer.
-Bleu: Maximum 15.17 from GPT-4o vs minimum 9.36 from Qwen2.5-7b, reflecting lexical alignment differences.
-5.5 Task III: Classification Experimentations
-In the depression risk binary classification task, we used three datasets, including the original SWDD dataset Cai et al.
-(2023) and the WU3D dataset Wang et al.
-(2020).
-Both of these datasets were randomly sampled with 200 negative users and 200 positive users for the experiment.
-Details provided in the table 4.
-In the classification experiments, in addition to the generative models mentioned in the section A.1, we also used models based on the BERT architecture with the details on A.7.
-6 Conclusion
-We introduced CNSocialDepress, a new dataset for depression risk analysis using Chinese social media texts.
-Uniquely, it is the first dataset to include both user-level and text-level labels, combining binary depression indicators with fine-grained six-dimensional analysis annotated by mental health professionals.
-We also developed an automated data generation pipeline to ease manual annotation and support language model fine-tuning for classification and summarization tasks.
-We believe CNSocialDepress will prove valuable across a range of applications, advancing early detection and intervention strategies in mental health.
-7 Limitations
-While CNSocialDepress provides a valuable resource for depression risk analyses in Chinese social media, several limitations should be noted.
-First, the dataset is sourced exclusively from Weibo, potentially introducing platform-specific demographic biases (e.g., underrepresentation of rural or elderly populations).
-Second, linguistic diversity, including dialectal variations and metaphorical expressions common in Chinese depressive discourse, remains insufficiently captured and may limit cross-regional generalizability.
-Third, although manual expert annotation ensures high quality, it constrains scalability, resulting in a relatively limited dataset size.
-Additionally, the current structured analysis focus primarily on depressive symptoms, omitting comorbid mental health conditions (e.g., anxiety) and contextual social factors.
-Future work will expand data collection to multi-platform sources, integrate dynamic lexicon updates for emerging expressions, and adopt hybrid annotation frameworks (e.g., expert-guided crowdsourcing) to enhance coverage and efficiency.
-Privacy-preserving techniques and longitudinal data tracking will further strengthen ethical and practical utility.
-8 Ethic Statement
-This dataset is built upon the SWDD Cai et al.
-(2023) benchmark released in early 2023.
-The original data comes from user-level text on the Weibo platform and was annotated by a team of psychology experts according to the DSM-5 symptom standards.
-During our work on the re-annotation process, we applied a second round of deep anonymization to meet privacy requirements.
-This involved systematically removing personal information, geographic locations, and potential identifiers from the original text, ensuring compliance with data protection protocols Benton et al.
-(2017) and GDPR (General Data Protection Regulation [^8] EUR-Lex GDPR Regulation).
-We also invited experienced researchers specializing in depression-related psychology to provide professional annotations based on their expertise.
-This dataset is mainly used for building early-warning models for depression risk.
-Through multi-dimensional feature analysis, it detects potential risk signals and can be used to automatically generate structured psychological assessment reports, offering technical support for digital mental health services.
-It is important to note that this study is strictly limited to the realm of mental health assistance, and any results should only serve as references for professional diagnosis or as a digital tool for self-assessment.
-At present, none of the algorithmic models can replace in-person psychiatric evaluations or deliver pathological diagnoses.
-If a system built using this dataset flags high-risk cases, or if individuals already show persistent mood disturbances or impaired social functioning, they should seek professional medical attention promptly.
-Mental health issues require clinical expertise, and any online self-assessment tool cannot substitute for in-person medical diagnosis.
+Table 6: Model performance on the depression risk summarization and analysis generation task.
+$5.$  Task III: Classification Experiments
+Our dataset is primarily designed for generation tasks.
+We also conduct a binary classification experiment for depression risk.
+In addition to our dataset, we use the original SWDD and WU3D datasets.
+For each dataset, we randomly sample 200 positive and 200 negative users.
+Besides the generative models described in Appendix A.1, we evaluate BERT-based classifiers (Devlin et al., 2019).
+Implementation details are provided in Appendix A.5.
+Results are reported in Appendix A.6.
+6.  Conclusion
+We introduce CNSocialDepress, a new dataset for depression risk analysis from Chinese social media.
+To our knowledge, it is among the first to provide both user-level and text-level labels.
+The dataset combines binary depression indicators with fine-grained six-dimensional analyses annotated by mental health professionals.
+We also develop an automated data generation pipeline.
+It reduces manual annotation effort and supports language model fine-tuning for classification and summarization.
+We expect CNSocialDepress to be useful for a range of research and applications, supporting early risk identification and downstream mental health interventions.
+7.  Limitations
+While CNSocialDepress provides a valuable resource for depression risk analysis on Chinese social media, several limitations should be noted.
+First, the dataset is sourced exclusively from Weibo.
+This may introduce platform-specific demographic biases, such as the underrepresentation of rural or elderly populations.
+Second, linguistic diversity is not fully captured.
+This includes dialectal variation and metaphorical expressions that are common in Chinese depressive discourse.
+This limitation may reduce generalizability across regions.
+Third, expert annotation ensures high quality, but it limits scalability.
+It also results in a relatively small dataset.
+In addition, the current structured analyses focus primarily on depressive symptoms.
+They omit comorbid mental health conditions (e.g., anxiety) and broader social context.
+Future work will expand data collection to multiple platforms.
+We also plan to update lexicons to capture emerging expressions.
+Finally, we will explore hybrid annotation frameworks (e.g., expert-guided crowdsourcing) to improve coverage and efficiency.
+Privacy-preserving techniques and longitudinal tracking can further strengthen ethical and practical utility.
+8.  Ethics Statement
+This dataset is built upon the SWDD benchmark Cai et al.
+(2023), released in early 2023.
+The original data consist of user-level posts from Weibo.
+They were annotated by psychology experts based on DSM-5 criteria.
+During re-annotation, we applied a second round of de-identification to protect privacy.
+We remove personal information, geographic locations, and other potential identifiers from the texts.
+We follow recommended data protection practices Benton et al.
+(2017) and comply with the GDPR (General Data Protection Regulation).
+[^7] https://eur-lex.europa.eu/legal-content/EN/TXT/PDF/?uri=CELEX:02016R0679-20160504 We also recruit experienced researchers in depression-related psychology to provide expert annotations.
+This dataset is intended for research on depression risk identification and for building assistive models.
+It can support multi-dimensional analysis of risk-related signals and the generation of structured summaries.
+It is not designed for clinical diagnosis, triage, or treatment decisions.
+Any outputs should be used only as references for professionals or as a self-assessment aid.
+No algorithmic system can replace in-person psychiatric evaluation or provide a definitive diagnosis.
+If a system built on this dataset flags a high-risk case, users should be encouraged to seek professional help.
+The same applies to individuals with persistent mood disturbances or impaired functioning.
+Online self-assessment tools cannot substitute for clinical evaluation.
+9.  Bibliographical References
+[…]
+  • Aldarwish and Ahmad (2017) Maryam Mohammed Aldarwish and Hafiz Farooq Ahmad.
+2017.
+Predicting depression levels using social media posts.
+In 2017 IEEE 13th international Symposium on Autonomous decentralized system (ISADS), pages 277–280.
+IEEE.
+  • Aleem et al.
+(2022) Shumaila Aleem, Noor ul Huda, Rashid Amin, Samina Khalid, Sultan S Alshamrani, and Abdullah Alshehri.
+2022.
+Machine learning algorithms for depression: diagnosis, insights, and research directions.
+Electronics, 11(7):1111.
+  • Alhamed et al.
+(2024) Falwah Alhamed, Julia Ive, and Lucia Specia.
+2024.
+Classifying social media users before and after depression diagnosis via their language usage: A dataset and study.
+In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 3250–3260, Torino, Italia.
+ELRA and ICCL.
+  • Almouzini et al. (2019) Salma Almouzini, Asem Alageel, et al. 2019. Detecting arabic depressed users from twitter data. Procedia Computer Science, 163:257–265.
+  • Amanat et al.
+(2022) Amna Amanat, Muhammad Rizwan, Abdul Rehman Javed, Maha Abdelhaq, Raed Alsaqour, Sharnil Pandya, and Mueen Uddin.
+2022.
+Deep learning for depression detection from textual data.
+Electronics, 11(5):676.
+  • Arora and Arora (2019) Priyanka Arora and Parul Arora.
+2019.
+Mining twitter data for depression detection.
+In 2019 international conference on signal processing and communication (ICSC), pages 186–189.
+IEEE.
+  • Association (2013) American Psychiatric Association. 2013. Diagnostic and statistical manual of mental disorders. Am Psychiatric Assoc, 21(21):591–643.
+  • Bai et al. (2023) Jinze Bai, Shuai Bai, Yunfei Chu, Zeyu Cui, Kai Dang, Xiaodong Deng, Yang Fan, Wenbin Ge, Yu Han, Fei Huang, et al. 2023. Qwen technical report. arXiv preprint arXiv:2309.16609.
+  • Bathina et al.
+(2021) Krishna C Bathina, Marijn Ten Thij, Lorenzo Lorenzo-Luaces, Lauren A Rutter, and Johan Bollen.
+2021.
+Individuals with depression express more distorted thinking on social media.
+Nature human behaviour, 5(4):458–466.
+  • Beck et al.
+(1996) Aaron T Beck, Robert A Steer, Roberta Ball, and William F Ranieri.
+1996.
+Comparison of beck depression inventories-ia and-ii in psychiatric outpatients.
+Journal of personality assessment, 67(3):588–597.
+  • Benton et al.
+(2017) Adrian Benton, Glen Coppersmith, and Mark Dredze.
+2017.
+Ethical research protocols for social media health research.
+In Proceedings of the First ACL Workshop on Ethics in Natural Language Processing, pages 94–102, Valencia, Spain.
+Association for Computational Linguistics.
+  • Bittar et al.
+(2019) André Bittar, Sumithra Velupillai, Angus Roberts, and Rina Dutta.
+2019.
+Text classification to inform suicide risk assessment in electronic health records.
+In MEDINFO 2019: Health and Wellbeing e-Networks for All, pages 40–44.
+IOS Press.
+  • Breiman (2001) Leo Breiman. 2001. Random forests. Machine learning, 45:5–32.
+  • Brown et al.
+(1992) Peter F Brown, Vincent J Della Pietra, Peter V Desouza, Jennifer C Lai, and Robert L Mercer.
+1992.
+Class-based n-gram models of natural language.
+Computational linguistics, 18(4):467–480.
+  • Bucur et al.
+(2025) Ana-Maria Bucur, Andreea Moldovan, Krutika Parvatikar, Marcos Zampieri, Ashiqur Khudabukhsh, and Liviu Dinu.
+2025.
+Datasets for depression modeling in social media: An overview.
+In Proceedings of the 10th Workshop on Computational Linguistics and Clinical Psychology (CLPsych 2025), pages 116–126, Albuquerque, New Mexico.
+Association for Computational Linguistics.
+  • Cacheda et al.
+(2019) Fidel Cacheda, Diego Fernandez, Francisco J Novoa, and Victor Carneiro.
+2019.
+Early detection of depression: social network analysis and random forest techniques.
+Journal of medical Internet research, 21(6):e12554.
+  • Cai et al.
+(2025a) Chengkun Cai, Haoliang Liu, Xu Zhao, Zhongyu Jiang, Tianfang Zhang, Zongkai Wu, John Lee, Jenq-Neng Hwang, and Lei Li.
+2025a.
+Bayesian optimization for controlled image editing via llms.
+In Proceedings of the Annual Meeting of the Association for Computational Linguistics (ACL).
+  • Cai et al.
+(2025b) Chengkun Cai, Xu Zhao, Haoliang Liu, Zhongyu Jiang, Tianfang Zhang, Zongkai Wu, Jenq-Neng Hwang, and Lei Li.
+2025b.
+The role of deductive and inductive reasoning in large language models.
+In Proceedings of the Annual Meeting of the Association for Computational Linguistics (ACL).
+  • Cai et al.
+(2023) Yicheng Cai, Haizhou Wang, Huali Ye, Yanwen Jin, and Wei Gao.
+2023.
+Depression detection on online social network with multivariate time series feature of user depressive symptoms.
+Expert Systems with Applications, 217:119538.
+  • Cha et al.
+(2022) Junyeop Cha, Seoyun Kim, and Eunil Park.
+2022.
+A lexicon-based approach to examine depression detection in social media: the case of twitter and university community.
+Humanities and Social Sciences Communications, 9(1):1–10.
+  • Chancellor and De Choudhury (2020) Stevie Chancellor and Munmun De Choudhury.
+2020.
+Methods in predictive techniques for mental health status on social media: a critical review.
+NPJ digital medicine, 3(1):43.
+  • Cortes (1995) Corinna Cortes. 1995. Support-vector networks. Machine Learning.
+  • Cui et al.
+(2020) Yiming Cui, Wanxiang Che, Ting Liu, Bing Qin, Shijin Wang, and Guoping Hu.
+2020.
+Revisiting pre-trained models for Chinese natural language processing.
+In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: Findings, pages 657–668, Online.
+Association for Computational Linguistics.
+  • Cui et al.
+(2019) Yiming Cui, Wanxiang Che, Ting Liu, Bing Qin, Ziqing Yang, Shijin Wang, and Guoping Hu.
+2019.
+Pre-training with whole word masking for chinese bert.
+arXiv preprint arXiv:1906.08101.
+  • DeepSeek-AI et al.
+(2025) DeepSeek-AI, Daya Guo, Dejian Yang, Haowei Zhang, Junxiao Song, Ruoyu Zhang, Runxin Xu, Qihao Zhu, Shirong Ma, Peiyi Wang, Xiao Bi, Xiaokang Zhang, Xingkai Yu, Yu Wu, Z.
+F.
+Wu, Zhibin Gou, Zhihong Shao, Zhuoshu Li, Ziyi Gao, Aixin Liu, et al.
+2025.
+Deepseek-r1: Incentivizing reasoning capability in llms via reinforcement learning.
+  • Devlin et al.
+(2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
+2019.
+BERT: Pre-training of deep bidirectional transformers for language understanding.
+In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 4171–4186, Minneapolis, Minnesota.
+Association for Computational Linguistics.
+  • Dreiseitl and Ohno-Machado (2002) Stephan Dreiseitl and Lucila Ohno-Machado.
+2002.
+Logistic regression and artificial neural network classification models: a methodology review.
+Journal of biomedical informatics, 35(5-6):352–359.
+  • Du et al.
+(2022) Zhiyuan Du, Yuxian Qian, Xiao Liu, Ming Ding, Yuxian Qian, Xiaoyang Liu, Jiezhong Qiu, Zhilin Yang, and Jie Tang.
+2022.
+Glm: General language model pretraining with autoregressive blank infilling.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 320–335.
+Association for Computational Linguistics.
+  • Eensoo and Valette (2012) Egle Eensoo and Mathieu Valette.
+2012.
+Sur l’application de méthodes textométriques à la construction de critères de classification en analyse des sentiments.
+In TALN 2012, volume 2, pages 367–374.
+GETALP-LIG.
+  • Evans-Lacko et al.
+(2018) SAGS Evans-Lacko, Sergio Aguilar-Gaxiola, Ali Al-Hamzawi, Jordi Alonso, Corina Benjet, Ronny Bruffaerts, Wai-Tat Chiu, Silvia Florescu, Giovanni de Girolamo, Oye Gureje, et al.
+2018.
+Socio-economic variations in the mental health treatment gap for people with anxiety, mood, and substance use disorders: results from the who world mental health (wmh) surveys.
+Psychological medicine, 48(9):1560–1571.
+  • Fernandes et al.
+(2018) Andrea C Fernandes, Rina Dutta, Sumithra Velupillai, Jyoti Sanyal, Robert Stewart, and David Chandran.
+2018.
+Identifying suicide ideation and suicidal attempts in a psychiatric clinical research database using natural language processing.
+Scientific reports, 8(1):7426.
+  • Gao et al.
+(2021) Shang Gao, Mohammed Alawad, M Todd Young, John Gounley, Noah Schaefferkoetter, Hong Jun Yoon, Xiao-Cheng Wu, Eric B Durbin, Jennifer Doherty, Antoinette Stroup, et al.
+2021.
+Limitations of transformers on clinical text classification.
+IEEE journal of biomedical and health informatics, 25(9):3596–3607.
+  • GeminiTeam et al.
+(2023) GeminiTeam, Rohan Anil, Sebastian Borgeaud, Yonghui Wu, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, et al.
+2023.
+Gemini: a family of highly capable multimodal models.
+arXiv preprint arXiv:2312.11805.
+  • GLMTeam et al.
+(2024) GLMTeam, Aohan Zeng, Bin Xu, Bowen Wang, Chenhui Zhang, Da Yin, Diego Rojas, Guanyu Feng, Hanlin Zhao, Hanyu Lai, Hao Yu, Hongning Wang, Jiadai Sun, Jiajie Zhang, Jiale Cheng, Jiayi Gui, Jie Tang, Jing Zhang, Juanzi Li, Lei Zhao, et al.
+2024.
+Chatglm: A family of large language models from glm-130b to glm-4 all tools.
+  • Graves (2013) Alex Graves. 2013. Generating sequences with recurrent neural networks. arXiv preprint arXiv:1308.0850.
+  • Gu et al.
+(2025) Ruocheng Gu, Sen Jia, Yule Ma, Jinqin Zhong, Jenq-Neng Hwang, and Lei Li.
+2025.
+Mocount: Motion-based repetitive action counting.
+In Proceedings of the 33rd ACM International Conference on Multimedia, pages 9026–9034.
+  • Guan et al.
+(2025) Yunchuan Guan, Yu Liu, Ke Zhou, Zhiqi Shen, Serge Belongie, Jenq-Neng Hwang, and Lei Li.
+2025.
+Learning to learn weight generation via local consistency diffusion.
+arXiv preprint arXiv:2502.01117.
+Accepted to CVPR 2026.
+  • Guntuku et al.
+(2017) Sharath Chandra Guntuku, David B Yaden, Margaret L Kern, Lyle H Ungar, and Johannes C Eichstaedt.
+2017.
+Detecting depression and mental illness on social media: an integrative review.
+Current Opinion in Behavioral Sciences, 18:43–49.
+  • Guo et al.
+(2023) Zhihua Guo, Nengneng Ding, Minyu Zhai, Zhenwen Zhang, and Zepeng Li.
+2023.
+Leveraging domain knowledge to improve depression detection on chinese social media.
+IEEE Transactions on Computational Social Systems, 10(4):1528–1536.
+  • Hämäläinen et al.
+(2021) Mika Hämäläinen, Pattama Patpong, Khalid Alnajjar, Niko Partanen, and Jack Rueter.
+2021.
+Detecting depression in Thai blog posts: a dataset and a baseline.
+In Proceedings of the Seventh Workshop on Noisy User-generated Text (W-NUT 2021), pages 20–25, Online.
+Association for Computational Linguistics.
+  • Harrigian et al.
+(2021) Keith Harrigian, Carlos Aguirre, and Mark Dredze.
+2021.
+On the state of social media data for mental health research.
+In Proceedings of the Seventh Workshop on Computational Linguistics and Clinical Psychology: Improving Access, pages 15–24, Online.
+Association for Computational Linguistics.
+  • Hasib et al.
+(2023) Khan Md Hasib, Md Rafiqul Islam, Shadman Sakib, Md Ali Akbar, Imran Razzak, and Mohammad Shafiul Alam.
+2023.
+Depression detection from social networks data based on machine learning and deep learning techniques: An interrogative survey.
+IEEE Transactions on Computational Social Systems, 10(4):1568–1586.
+  • He and Qu (2025a) Qi He and Chunyu Qu. 2025a. Modular landfill remediation for ai grid resilience. arXiv preprint.
+  • He and Qu (2025b) Qi He and Chunyu Qu. 2025b. Waste-to-energy-coupled ai data centers: Cooling efficiency and grid resilience. arXiv preprint.
+  • He et al.
+(2026) Qi He, Rui Shan, Chunyu Qu, Yuan Tang, and Yue Zou.
+2026.
+Rare-earth exposure and bottlenecks in ai data centers: Cost and schedule risk.
+SSRN Electronic Journal.
+Available at SSRN.
+  • Hu et al.
+(2022) Edward J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen.
+2022.
+LoRA: Low-rank adaptation of large language models.
+In International Conference on Learning Representations.
+  • Hu et al. (2024) Jinpeng Hu, Tengteng Dong, Luo Gang, Hui Ma, Peng Zou, Xiao Sun, Dan Guo, and Meng Wang. 2024. Psycollm: Enhancing llm for psychological understanding and evaluation.
+  • Islam et al.
+(2018) Md Rafiqul Islam, Muhammad Ashad Kabir, Ashir Ahmed, Abu Raihan M Kamal, Hua Wang, and Anwaar Ulhaq.
+2018.
+Depression detection from social network data using machine learning techniques.
+Health information science and systems, 6:1–12.
+  • Ji et al.
+(2022) Shaoxiong Ji, Tianyu Zhang, Laith Ansari, Jie Fu, Piyush Tiwari, and Erik Cambria.
+2022.
+Mentalbert: Publicly available pretrained language models for mental healthcare.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 7184–7190.
+European Language Resources Association.
+  • Jia and Li (2024) Sen Jia and Lei Li. 2024. Adaptive masking enhances visual grounding. arXiv preprint arXiv:2410.03161.
+  • Jiang et al.
+(2025) Zhongyu Jiang, Wenhao Chai, Lei Li, Zhuoran Zhou, Cheng-Yen Yang, and Jenq-Neng Hwang.
+2025.
+Unihpr: Unified human pose representation via singular value contrastive learning.
+arXiv preprint arXiv:2510.19078.
+  • Jiang et al.
+(2024) Zhongyu Jiang, Zhuoran Zhou, Lei Li, Wenhao Chai, Cheng-Yen Yang, and Jenq-Neng Hwang.
+2024.
+Back to optimization: Diffusion-based zero-shot 3d human pose estimation.
+In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision (WACV).
+  • Joulin et al.
+(2016) Armand Joulin, Edouard Grave, Piotr Bojanowski, Matthijs Douze, Hérve Jégou, and Tomas Mikolov.
+2016.
+Fasttext.zip: Compressing text classification models.
+arXiv preprint arXiv:1612.03651.
+  • Kaya et al.
+(2019) Heysem Kaya, Dmitrii Fedotov, Denis Dresvyanskiy, Metehan Doyran, Danila Mamontov, Maxim Markitantov, Alkim Almila Akdag Salah, Evrim Kavcar, Alexey Karpov, and Albert Ali Salah.
+2019.
+Predicting depression and emotions in the cross-roads of cultures, para-linguistics, and non-linguistics.
+In Proceedings of the 9th International on Audio/Visual Emotion Challenge and Workshop, pages 27–35.
+  • Kim (2014) Yoon Kim.
+2014.
+Convolutional neural networks for sentence classification.
+In Proceedings of the 2014 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1746–1751, Doha, Qatar.
+Association for Computational Linguistics.
+  • Kroenke et al. (2001) Kurt Kroenke, Robert L Spitzer, and Janet BW Williams. 2001. The phq-9: validity of a brief depression severity measure. Journal of general internal medicine, 16(9):606–613.
+  • Lafon (1980) Pierre Lafon. 1980. Sur la variabilité de la fréquence des formes dans un corpus. Mots. Les langages du politique, 1(1):127–165.
+  • Lai et al.
+(2023) Tin Lai, Yukun Shi, Zicong Du, Jiajie Wu, Ken Fu, Yichao Dou, and Ziqi Wang.
+2023.
+Psy-llm: Scaling up global mental health psychological services with ai-based large language models.
+arXiv preprint arXiv:2307.11991.
+  • Lan et al. (2024) Xiaochong Lan, Yiming Cheng, Li Sheng, Chen Gao, and Yong Li. 2024. Depression detection on social media with large language models. arXiv preprint arXiv:2403.10750.
+  • Le Glaz et al.
+(2021) Aziliz Le Glaz, Yannis Haralambous, Deok-Hee Kim-Dufor, Philippe Lenca, Romain Billot, Taylor C Ryan, Jonathan Marsh, Jordan Devylder, Michel Walter, Sofian Berrouiguet, et al.
+2021.
+Machine learning and natural language processing in mental health: systematic review.
+Journal of medical Internet research, 23(5):e15708.
+  • Li et al.
+(2020) Genghao Li, Bing Li, Langlin Huang, Sibing Hou, et al.
+2020.
+Automatic construction of a depression-domain lexicon based on microblogs: text mining study.
+JMIR medical informatics, 8(6):e17650.
+  • Li et al.
+(2025a) Hao Li, Ju Dai, Xin Zhao, Feng Zhou, Junjun Pan, and Lei Li.
+2025a.
+Wav2sem: Plug-and-play audio semantic decoupling for 3d speech-driven facial animation.
+arXiv preprint arXiv:2505.23290.
+  • Li et al. (2025b) Hao Li, Ju Dai, Feng Zhou, Kaida Ning, Lei Li, and Junjun Pan. 2025b. Au-blendshape for fine-grained stylized 3d facial expression manipulation. arXiv preprint arXiv:2507.12001.
+  • Li et al. (2026) Lei Li, Sen Jia, and Jenq-Neng Hwang. 2026. Multiple human motion understanding. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 40, pages 6297–6305.
+  • Li et al.
+(2025c) Lei Li, Sen Jia, Jianhao Wang, Zhaochong An, Jiaang Li, Jenq-Neng Hwang, and Serge Belongie.
+2025c.
+Chatmotion: A multimodal multi-agent for human motion analysis.
+arXiv preprint arXiv:2502.18180.
+  • Li et al.
+(2025d) Lei Li, Sen Jia, Jianhao Wang, Zhongyu Jiang, Feng Zhou, Ju Dai, Tianfang Zhang, Zongkai Wu, and Jenq-Neng Hwang.
+2025d.
+Human Motion Instruction Tuning.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
+  • Li et al. (2022a) Lei Li, Tianfang Zhang, Zhongfeng Kang, and Xikun Jiang. 2022a. Mask-fpan: Semi-supervised face parsing in the wild with de-occlusion and uv gan. arXiv preprint arXiv:2212.09098.
+  • Li et al.
+(2022b) Mingzheng Li, Haojie Xu, Weifeng Liu, and Jiangwei Liu.
+2022b.
+Bidirectional lstm and attention for depression detection on clinical interview transcripts.
+In 2022 IEEE 10th International Conference on Information, Communication and Networks (ICICN), pages 638–643.
+IEEE.
+  • Li et al.
+(2023) Zepeng Li, Zhengyi An, Wenchuan Cheng, Jiawei Zhou, Fang Zheng, and Bin Hu.
+2023.
+Mha: a multimodal hierarchical attention model for depression detection in social media.
+Health information science and systems, 11(1):6.
+  • Lin et al.
+(2020) Chenhao Lin, Pengwei Hu, Hui Su, Shaochun Li, Jing Mei, Jie Zhou, and Henry Leung.
+2020.
+Sensemood: depression detection on social media.
+In Proceedings of the 2020 international conference on multimedia retrieval, pages 407–411.
+  • Lin (2004) Chin-Yew Lin. 2004. ROUGE: A package for automatic evaluation of summaries. In Text Summarization Branches Out, pages 74–81, Barcelona, Spain. Association for Computational Linguistics.
+  • Liu et al.
+(2024) Libin Liu, Shen Chen, Sen Jia, Jingzhe Shi, Zhongyu Jiang, Can Jin, Wu Zongkai, Jenq-Neng Hwang, and Lei Li.
+2024.
+Graph canvas for controllable 3d scene generation.
+arXiv preprint arXiv:2412.00091.
+  • Liu et al.
+(2019a) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
+2019a.
+Roberta: A robustly optimized bert pretraining approach.
+  • Liu et al.
+(2019b) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
+2019b.
+Roberta: A robustly optimized bert pretraining approach.
+arXiv preprint arXiv:1907.11692.
+  • Maghraby and Ali (2022) Ashwag Maghraby and Hosnia Ali. 2022. Modern standard arabic mood changing and depression dataset. Data in Brief, 41:107999.
+  • Malhotra and Jindal (2022) Anshu Malhotra and Rajni Jindal.
+2022.
+Deep learning techniques for suicide and depression detection from online social media: A scoping review.
+Applied Soft Computing, 130:109713.
+  • Malviya et al.
+(2021) Keshu Malviya, Bholanath Roy, and SK Saritha.
+2021.
+A transformers approach to detect depression in social media.
+In 2021 International Conference on Artificial Intelligence and Smart Systems (ICAIS), pages 718–723.
+IEEE.
+  • McCallum et al.
+(1998) Andrew McCallum, Kamal Nigam, et al.
+1998.
+A comparison of event models for naive bayes text classification.
+In AAAI-98 workshop on learning for text categorization, volume 752, pages 41–48.
+Madison, WI.
+  • Mikolov et al. (2013) Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013. Efficient estimation of word representations in vector space.
+  • Moiseev et al.
+(2022) Fedor Moiseev, Zhe Dong, Enrique Alfonseca, and Martin Jaggi.
+2022.
+SKILL: Structured knowledge infusion for large language models.
+In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 1581–1588, Seattle, United States.
+Association for Computational Linguistics.
+  • Mothe et al.
+(2022) Josiane Mothe, Faneva Ramiandrisoa, and Md Zia Ullah.
+2022.
+Comparison of machine learning models for early depression detection from users’ posts.
+In Early Detection of Mental Health Disorders by Social Media Monitoring: The First Five Years of the eRisk Project, volume 1018 of Studies in Computational Intelligence book series (SCI), pages 111–139.
+Springer International Publishing.
+  • Mowery et al.
+(2017) Danielle Mowery, Hilary Smith, Tyler Cheney, Greg Stoddard, Glen Coppersmith, Craig Bryan, Mike Conway, et al.
+2017.
+Understanding depressive symptoms and psychosocial stressors on twitter: a corpus-based study.
+Journal of medical Internet research, 19(2):e6895.
+  • Orabi et al.
+(2018) Ahmed Husseini Orabi, Prasadith Buddhitha, Mahmoud Husseini Orabi, and Diana Inkpen.
+2018.
+Deep learning for depression detection of twitter users.
+In Proceedings of the fifth workshop on computational linguistics and clinical psychology: from keyboard to clinic, pages 88–97.
+  • Papineni et al.
+(2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
+2002.
+Bleu: a method for automatic evaluation of machine translation.
+In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, pages 311–318, Philadelphia, Pennsylvania, USA.
+Association for Computational Linguistics.
+  • Parapar et al.
+(2022) Javier Parapar, Patricia Martín-Rodilla, David E.
+Losada, and Fabio Crestani.
+2022.
+Overview of erisk at clef 2022: Early risk prediction on the internet (extended overview).
+In CLEF 2022: Conference and Labs of the Evaluation Forum, September 5–8, 2022, Bologna, Italy, Bologna, Italy.
+CEUR Workshop Proceedings.
+CC BY 4.0; ISSN: 1613-0073.
+  • Pennington et al.
+(2014) Jeffrey Pennington, Richard Socher, and Christopher Manning.
+2014.
+GloVe: Global vectors for word representation.
+In Proceedings of the 2014 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1532–1543, Doha, Qatar.
+Association for Computational Linguistics.
+  • Pincemin (2022) Bénédicte Pincemin. 2022. Sémantique textométrique. La sémantique au pluriel. Théories et méthodes, pages 373–396.
+  • QwenTeam (2024) QwenTeam. 2024. Qwen2.5: A party of foundation models.
+  • Raihan et al.
+(2024) Nishat Raihan, Sadiya Sayara Chowdhury Puspo, Shafkat Farabi, Ana-Maria Bucur, Tharindu Ranasinghe, and Marcos Zampieri.
+2024.
+MentalHelp: A multi-task dataset for mental health in social media.
+In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 11196–11203, Torino, Italia.
+ELRA and ICCL.
+  • Ramos et al.
+(2003) Juan Ramos et al.
+2003.
+Using tf-idf to determine word relevance in document queries.
+In Proceedings of the first instructional conference on machine learning, volume 242, pages 29–48.
+Citeseer.
+  • Romero et al.
+(2024) Alba M Mármol Romero, Adrián Moreno Muñoz, Flor Miriam Plaza Del Arco, M Dolores Molina-González, María Teresa Martín Valdivia, L Alfonso Urena Lopez, and Arturo Montejo Ráez.
+2024.
+Mentalriskes: A new corpus for early detection of mental disorders in spanish.
+In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 11204–11214.
+  • Sak et al.
+(2014) Haşim Sak, Andrew Senior, and Françoise Beaufays.
+2014.
+Long short-term memory based recurrent neural network architectures for large vocabulary speech recognition.
+arXiv preprint arXiv:1402.1128.
+  • Salem (1986a) André Salem. 1986a. Segments répétés et analyse statistique des données textuelles. Histoire & Mesure, 1:5–28.
+  • Salem (1986b) André Salem. 1986b. Segments répétés et analyse statistique des données textuelles. Histoire & mesure, pages 5–28.
+  • Santos et al.
+(2024) Wesley Ramos dos Santos, Rafael Lage de Oliveira, and Ivandré Paraboni.
+2024.
+Setembrobr: a social media corpus for depression and anxiety disorder prediction.
+Language Resources and Evaluation, 58(1):273–300.
+  • Schmidhuber (2015) Jürgen Schmidhuber. 2015. Deep learning in neural networks: An overview. Neural Networks, 61:85–117.
+  • Seabrook et al.
+(2018) Elizabeth M Seabrook, Margaret L Kern, Ben D Fulcher, and Nikki S Rickard.
+2018.
+Predicting depression from language-based emotion dynamics: longitudinal analysis of facebook and twitter status updates.
+Journal of medical Internet research, 20(5):e168.
+  • Shah et al.
+(2020) Faisal Muhammad Shah, Farzad Ahmed, Sajib Kumar Saha Joy, Sifat Ahmed, Samir Sadek, Rimon Shil, and Md Hasanul Kabir.
+2020.
+Early depression detection from social network using deep learning techniques.
+In 2020 IEEE region 10 symposium (TENSYMP), pages 823–826.
+IEEE.
+  • Shen et al.
+(2017) Guangyao Shen, Jia Jia, Liqiang Nie, Fuli Feng, Cunjun Zhang, Tianrui Hu, Tat-Seng Chua, Wenwu Zhu, et al.
+2017.
+Depression detection via harvesting social media: A multimodal dictionary learning solution.
+In IJCAI, pages 3838–3844.
+  • Shen et al.
+(2018) Tiancheng Shen, Jia Jia, Guangyao Shen, Fuli Feng, Xiangnan He, Huanbo Luan, Jie Tang, Thanassis Tiropanis, Tat Seng Chua, and Wendy Hall.
+2018.
+Cross-domain depression detection via harvesting social media.
+In Proceedings of the International Joint Conference on Artificial Intelligence.
+International Joint Conferences on Artificial Intelligence.
+  • Shen et al.
+(2022) Ying Shen, Huiyu Yang, and Lin Lin.
+2022.
+Automatic depression detection: An emotional audio-textual corpus and a gru/bilstm-based model.
+In ICASSP 2022-2022 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), pages 6247–6251.
+IEEE.
+  • Shi et al.
+(2025) Jingzhe Shi, Qinwei Ma, Hongyi Liu, Hang Zhao, Jeng-Neng Hwang, and Lei Li.
+2025.
+Explaining context length scaling and bounds for language models.
+arXiv preprint arXiv:2502.01481.
+  • Shi et al. (2024) Jingzhe Shi, Qinwei Ma, Huan Ma, and Lei Li. 2024. Scaling law for time series forecasting. In Advances in Neural Information Processing Systems (NeurIPS).
+  • (104) Pengcheng Shi, Jiawei Chen, Jiaqi Liu, Xinglin Zhang, Tao Chen, and Lei Li.
+Medal s: Spatio-textual prompt model for medical segmentation.
+In CVPR 2025: Foundation Models for 3D Biomedical Image Segmentation.
+  • Sotudeh et al.
+(2021) Sajad Sotudeh, Hanieh Deilamsalehy, Franck Dernoncourt, and Nazli Goharian.
+2021.
+TLDR9+: A large scale resource for extreme summarization of social media posts.
+In Proceedings of the Third Workshop on New Frontiers in Summarization, pages 142–151, Online and in Dominican Republic.
+Association for Computational Linguistics.
+  • Sotudeh et al.
+(2022) Sajad Sotudeh, Nazli Goharian, and Zachary Young.
+2022.
+MentSum: A resource for exploring summarization of mental health online posts.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 2682–2692, Marseille, France.
+European Language Resources Association.
+  • Squires et al.
+(2023) Matthew Squires, Xiaohui Tao, Soman Elangovan, Raj Gururajan, Xujuan Zhou, U Rajendra Acharya, and Yuefeng Li.
+2023.
+Deep learning and machine learning in psychiatry: a survey of current progress in depression detection, diagnosis and treatment.
+Brain Informatics, 10(1):10.
+  • Stankevich et al.
+(2020) Maxim Stankevich, Ivan Smirnov, Natalia Kiselnikova, and Anastasia Ushakova.
+2020.
+Depression detection from social media profiles.
+In Data Analytics and Management in Data Intensive Domains: 21st International Conference, DAMDID/RCDL 2019, Kazan, Russia, October 15–18, 2019, Revised Selected Papers 21, pages 181–194.
+Springer.
+  • Tausczik and Pennebaker (2010) Yla R Tausczik and James W Pennebaker.
+2010.
+The psychological meaning of words: Liwc and computerized text analysis methods.
+Journal of language and social psychology, 29(1):24–54.
+  • Touvron et al.
+(2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al.
+2023.
+Llama: Open and efficient foundation language models.
+arXiv preprint arXiv:2302.13971.
+  • Trotzek et al.
+(2018a) Marcel Trotzek, Sven Koitka, and Christoph M Friedrich.
+2018a.
+Utilizing neural networks and linguistic metadata for early detection of depression indications in text sequences.
+IEEE Transactions on Knowledge and Data Engineering, 32(3):588–601.
+  • Trotzek et al.
+(2018b) Marcel Trotzek, Sven Koitka, and Christoph M Friedrich.
+2018b.
+Word embeddings and linguistic metadata at the clef 2018 tasks for early detection of depression and anorexia.
+In CLEF (working notes).
+  • Un Nisa and Muhammad (2021) Qamar Un Nisa and Rafi Muhammad.
+2021.
+Towards transfer learning using bert for early detection of self-harm of social media users.
+Proceedings of the Working Notes of CLEF, pages 21–4.
+  • Vaswani (2017) A Vaswani. 2017. Attention is all you need. Advances in Neural Information Processing Systems.
+  • Wang et al.
+(2024a) Jifei Wang, Zhenping Zhao, Jing Yang, Limin Wang, Mei Zhang, and Maigeng Zhou.
+2024a.
+The association between depression and all-cause, cause-specific mortality in the chinese population—china, 2010–2022.
+China CDC Weekly, 6(40):1022.
+  • Wang et al. (2024b) Shenzhi Wang, Yaowei Zheng, Guoyin Wang, Shiji Song, and Gao Huang. 2024b. Llama3-8b-chinese-chat (revision 6622a23).
+  • Wang et al.
+(2019) Wei Wang, Bin Bi, Ming Yan, Chen Wu, Zuyi Bao, Jiangnan Xia, Liwei Peng, and Luo Si.
+2019.
+Structbert: Incorporating language structures into pre-training for deep language understanding.
+arXiv preprint arXiv:1908.04577.
+  • Wang et al.
+(2016) Yequan Wang, Minlie Huang, Xiaoyan Zhu, and Li Zhao.
+2016.
+Attention-based lstm for aspect-level sentiment classification.
+In Proceedings of the 2016 conference on empirical methods in natural language processing, pages 606–615.
+  • Wang et al.
+(2020a) Yiding Wang, Zhenyi Wang, Chenghao Li, Yilin Zhang, and Haizhou Wang.
+2020a.
+A multimodal feature fusion-based method for individual depression detection on sina weibo.
+In 2020 IEEE 39th International Performance Computing and Communications Conference (IPCCC), pages 1–8.
+  • Wang et al.
+(2020b) Yiding Wang, Zhenyi Wang, Chenghao Li, Yilin Zhang, and Haizhou Wang.
+2020b.
+A multimodal feature fusion-based method for individual depression detection on sina weibo.
+In 2020 IEEE 39th International Performance Computing and Communications Conference (IPCCC), pages 1–8.
+  • Wang et al.
+(2024c) Yuxi Wang, Diana Inkpen, and Prasadith Kirinde Gamaarachchige.
+2024c.
+Explainable depression detection using large language models on social media data.
+In Proceedings of the 9th Workshop on Computational Linguistics and Clinical Psychology (CLPsych 2024), pages 108–126.
+  • Wei et al.
+(2023) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed Chi, Quoc Le, and Denny Zhou.
+2023.
+Chain-of-thought prompting elicits reasoning in large language models.
+  • Wolohan et al.
+(2018) JT Wolohan, Misato Hiraga, Atreyee Mukherjee, Zeeshan Ali Sayyed, and Matthew Millard.
+2018.
+Detecting linguistic traces of depression in topic-restricted text: Attending to self-stigmatized depression with nlp.
+In Proceedings of the first international workshop on language cognition and computational models, pages 11–21.
+  • Wu et al.
+(2020) Min Yen Wu, Chih-Ya Shen, En Tzu Wang, and Arbee LP Chen.
+2020.
+A deep architecture for depression detection using posting, behavior, and living environment data.
+Journal of Intelligent Information Systems, 54(2):225–244.
+  • Xin Yan (2023) Dong Xue* Xin Yan. 2023. Mindchat: Psychological large language model. https://github.com/X-D-Lab/MindChat.
+  • Xu et al.
+(2025) Jinyuan Xu, Tian Lan, Mathieu Valette, Pierre Magistry, and Lei Li.
+2025.
+TinyMentalLLMs enable depression detection in Chinese social media texts.
+In Proceedings of the 15th International Conference on Recent Advances in Natural Language Processing - Natural Language Processing in the Generative AI Era, pages 1352–1363, Varna, Bulgaria.
+INCOMA Ltd., Shoumen, Bulgaria.
+  • Xu et al.
+(2024) Xuhai Xu, Bingsheng Yao, Yuanzhe Dong, Saadia Gabriel, Hong Yu, James Hendler, Marzyeh Ghassemi, Anind K Dey, and Dakuo Wang.
+2024.
+Mental-llm: Leveraging large language models for mental health prediction via online text data.
+Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies, 8(1):1–32.
+  • Yang et al.
+(2024a) An Yang, Baosong Yang, Binyuan Hui, Bo Zheng, Bowen Yu, Chang Zhou, Chengpeng Li, Chengyuan Li, Dayiheng Liu, Fei Huang, Guanting Dong, Haoran Wei, Huan Lin, Jialong Tang, Jialin Wang, Jian Yang, Jianhong Tu, Jianwei Zhang, Jianxin Ma, Jin Xu, et al.
+2024a.
+Qwen2 technical report.
+arXiv preprint arXiv:2407.10671.
+  • Yang and Li (2009) Juhua Yang and Lulu Li. 2009. Intergenerational dynamics and family solidarity: A comparative study of mainland china, japan, korea and taiwan. Sociological Studies, 3:26–53.
+  • Yang et al.
+(2023) Kailai Yang, Shaoxiong Ji, Tianlin Zhang, Qianqian Xie, Ziyan Kuang, and Sophia Ananiadou.
+2023.
+Towards interpretable mental health analysis with large language models.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 6056–6077, Singapore.
+Association for Computational Linguistics.
+  • Yang et al.
+(2024b) Kailai Yang, Tianlin Zhang, Ziyan Kuang, Qianqian Xie, Jimin Huang, and Sophia Ananiadou.
+2024b.
+Mentallama: interpretable mental health analysis on social media with large language models.
+In Proceedings of the ACM on Web Conference 2024, pages 4489–4500.
+  • Yang et al.
+(2021) Tingting Yang, Fei Li, Donghong Ji, Xiaohui Liang, Tian Xie, Shuwan Tian, Bobo Li, and Peitong Liang.
+2021.
+Fine-grained depression analysis based on chinese micro-blog reviews.
+Information Processing & Management, 58(6):102681.
+  • Yang et al. (2025) Wenhao Yang, Jianguo Wei, Wenhuan Lu, and Lei Li. 2025. You only speak once to see. In IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP).
+  • Yang et al.
+(2020) Xingwei Yang, Rhonda McEwen, Liza Robee Ong, and Morteza Zihayat.
+2020.
+A big data analytics framework for detecting user-level depression from social networks.
+International Journal of Information Management, 54:102141.
+  • Yao et al.
+(2024) Yifan Yao, Jinhao Duan, Kaidi Xu, Yuanfang Cai, Zhibo Sun, and Yue Zhang.
+2024.
+A survey on large language model (llm) security and privacy: The good, the bad, and the ugly.
+High-Confidence Computing, page 100211.
+  • Yao et al.
+(2025) Ziyu Yao, Xuxin Cheng, Zhiqi Huang, and Lei Li.
+2025.
+CountLLM: Towards Generalizable Repetitive Action Counting via Large Language Model.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
+  • Yates et al.
+(2017) Andrew Yates, Arman Cohan, and Nazli Goharian.
+2017.
+Depression and self-harm risk assessment in online forums.
+In Proceedings of the 2017 Conference on Empirical Methods in Natural Language Processing, pages 2968–2978, Copenhagen, Denmark.
+Association for Computational Linguistics.
+  • Yazdavar et al.
+(2020) Amir Hossein Yazdavar, Mohammad Saeid Mahdavinejad, Goonmeet Bajaj, William Romine, Amit Sheth, Amir Hassan Monadjemi, Krishnaprasad Thirunarayan, John M Meddar, Annie Myers, Jyotishman Pathak, et al.
+2020.
+Multimodal mental health analysis in social media.
+Plos one, 15(4):e0226248.
+  • Ye et al. (2023) Hongbin Ye, Tong Liu, Aijia Zhang, Wei Hua, and Weiqiang Jia. 2023. Cognitive mirage: A review of hallucinations in large language models.
+  • Yuka Niimi (2021) Yutaka Miyaji Yuka Niimi.
+2021.
+Machine learning approach for depression detection in Japanese.
+In Proceedings of the 35th Pacific Asia Conference on Language, Information and Computation, pages 346–353, Shanghai, China.
+Association for Computational Linguistics.
+  • Zeiler and Fergus (2014) Matthew D Zeiler and Rob Fergus.
+2014.
+Visualizing and understanding convolutional networks.
+In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part I 13, pages 818–833.
+Springer.
+  • Zhai et al.
+(2024) Wei Zhai, Hongzhi Qi, Qing Zhao, Jianqiang Li, Ziqi Wang, Han Wang, Bing Yang, and Guanghui Fu.
+2024.
+Chinese MentalBERT: Domain-adaptive pre-training on social media for Chinese mental health text analysis.
+In Findings of the Association for Computational Linguistics ACL 2024, pages 10574–10585, Bangkok, Thailand and virtual meeting.
+Association for Computational Linguistics.
+  • Zhang et al.
+(2020) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q.
+Weinberger, and Yoav Artzi.
+2020.
+Bertscore: Evaluating text generation with BERT.
+In 8th International Conference on Learning Representations, ICLR 2020, Addis Ababa, Ethiopia, April 26-30, 2020.
+OpenReview.net.
+  • Zhang and Wallace (2015) Ye Zhang and Byron Wallace.
+2015.
+A sensitivity analysis of (and practitioners’ guide to) convolutional neural networks for sentence classification.
+arXiv preprint arXiv:1510.03820.
+  • Zhang et al.
+(2021) Yipeng Zhang, Hanjia Lyu, Yubao Liu, Xiyang Zhang, Yu Wang, and Jiebo Luo.
+2021.
+Monitoring depression trends on twitter during the covid-19 pandemic: observational study.
+JMIR infodemiology, 1(1):e26769.
+  • Zhang et al.
+(2022) Zhiling Zhang, Siyuan Chen, Mengyue Wu, and Kenny Zhu.
+2022.
+Symptom identification for interpretable detection of multiple mental disorders on social media.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 9970–9985, Abu Dhabi, United Arab Emirates.
+Association for Computational Linguistics.
+  • Zheng et al.
+(2024) Yaowei Zheng, Richong Zhang, Junhao Zhang, Yanhan Ye, Zheyan Luo, Zhangchi Feng, and Yongqiang Ma.
+2024.
+Llamafactory: Unified efficient fine-tuning of 100+ language models.
+In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations), Bangkok, Thailand.
+Association for Computational Linguistics.
+  • Zhou et al.
+(2023) Han-yu Zhou, Wen-qi Zhu, Wen-yi Xiao, Ya-ting Huang, Kang Ju, Hong Zheng, and Chao Yan.
+2023.
+Feeling unloved is the most robust sign of adolescent depression linking to family communication patterns.
+Journal of Research on Adolescence, 33(2):418–430.
+  • Zhou et al.
+(2024) Zhuoran Zhou, Zhongyu Jiang, Wenhao Chai, Cheng-Yen Yang, Lei Li, and Jenq-Neng Hwang.
+2024.
+Efficient domain adaptation via generative prior for 3d infant pose estimation.
+In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision (WACV).
 Appendix A Appendix
-A.1 Baseline Models
-  • DeepSeek-R1 is a large-scale model specialized in logical reasoning,This model excels in complex reasoning tasks such as mathematics and programming while maintaining strong overall performance on general tasks.
-Chosen for its specialized excellence in complex logical reasoning while maintaining strong general task performance, potentially beneficial for in-depth structured psychological analysis and nuanced depression risk detection.
-  • DeepSeek R1 Distill-14B is a lightweight reasoning model derived from Qwen2.5-14B through Knowledge Distillation.
-Chosen for its lightweight reasoning capabilities, aiming to retain the advantages of larger models in complex tasks while reducing computational cost, making it suitable for efficient depression risk analysis.
-  • Qwen2.5-14B is part of Alibaba’s Qwen series of large language models.
-Chosen for its strong performance in long-form text generation and instruction-following, relevant for generating structured psychological analyses from Chinese social media posts.
+A.1.  Baseline Models
+  • DeepSeek-R1 is a large-scale model designed for complex reasoning.
+It performs well on tasks such as mathematics and programming, while remaining competitive on general-purpose benchmarks.
+We include it as a strong reasoning-oriented baseline for structured psychological analysis and depression risk assessment.
+  • DeepSeek-R1-Distill-14B is a distilled reasoning model derived from Qwen2.5-14B via knowledge distillation.
+It aims to retain key reasoning capabilities at lower computational cost.
+We include it as an efficient baseline for depression risk analysis.
+  • Qwen2.5-14B is a large language model from the Qwen series.
+It provides strong instruction following and long-form generation.
+We include it as a Chinese-capable baseline for generating structured psychological analyses from social media posts.
   • GPT-4o is a multimodal large language model released by OpenAI.
-Chosen for its state-of-the-art multimodal and multilingual processing capabilities, including Chinese, and its improved overall performance and response speed, making it a powerful baseline for depression detection and psychological analysis.
-  • GPT-4o-Mini is a lightweight AI model developed by OpenAI to provide an efficient and cost-effective AI solution.
-Chosen as a resource- and cost-efficient model with robust text understanding and generation, suitable for less computationally demanding depression risk identification and basic psychological analysis tasks.
-  • Llama3-8B-Chinese-Chat, derived from Meta-Llama-3-8B, is an instruction-tuned model optimized for Chinese dialogue tasks.
-It has been fine-tuned to process both Chinese and English inputs.
-Chosen for its optimization for Chinese dialogue tasks and ability to handle extended conversations, making it well-suited for analyzing conversational social media posts for depression risk and generating coherent psychological insights in Chinese.
-A.2 20 Randomly Generated Expression for the Positive Users
+It supports multilingual generation, including Chinese.
+We include it as a strong general-purpose baseline for depression detection and psychological analysis.
+  • GPT-4o-mini is a smaller model in the GPT-4o family.
+It is designed for efficiency and lower inference cost.
+We include it as a lightweight baseline for depression risk identification and structured summarization.
+  • Llama3-8B-Chinese-Chat, derived from Meta-Llama-3-8B, is an instruction-tuned model for Chinese dialogue.
+It supports both Chinese and English inputs.
+We include it as a Chinese dialogue-oriented baseline for user-level depression risk analysis and generation.
+A.2.  20 Randomly Generated Expression for the Positive Users
   1. 1.
@@
     Note: Only indices and explanations, not original text.
-A.3 20 Distinct Label Expressions for the Negative Users
+A.3.  20 Distinct Label Expressions for the Negative Users
   1. 1.
@@
     "The content lacks descriptions of negative emotions and is therefore not classified under any depression dimensions."
-A.4 Instructions for the Fine-Tuning Process
+A.4.  Instructions for the Fine-Tuning Process
   • Complete Instruction = Task Instruction + Supplementary Instruction
@@
     "Please read the following text and determine whether it contains any one of the above 6 depressive emotion expression dimensions, and provide a concise explanation; if not, please indicate that the text does not belong to any depression dimension."
-A.5 Module II Prompt
+A.5.  BERT-based Models
+BERT-base-chinese [^8] https://huggingface.co/google-bert/bert-base-chinese follows the BERT-Base architecture Devlin et al.
+(2019) and is pre-trained on large-scale Chinese corpora.
+It uses masked language modeling (MLM) and next sentence prediction (NSP).
+The model learns bidirectional representations for Chinese text, typically at the character level.
+Chinese-RoBERTa-wwm-ext Cui et al.
+(2019) is based on RoBERTa Liu et al.
+(2019b) and uses whole word masking (WWM).
+It masks complete words rather than individual characters during MLM.
+This design can better capture word-level semantics in Chinese.
+The extended pre-training (“ext”) on large Chinese corpora further improves robustness across NLP tasks.
+StructBERT-mental Wang et al.
+(2019) builds on StructBERT, which incorporates structural objectives during pre-training.
+It emphasizes word- and sentence-level structure.
+In addition, it is adapted for mental health text, which can benefit mental health-related classification.
+A.6.  Classification Performance Comparison
+Model/Test Dataset Swdd Origin Wu3d CNSD (Gold)
+Acc F1 Acc F1 Acc F1
+RoBERTa Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.88 0.87
+RoBERTa Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.82 0.78
+Bert based Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.85 0.82
+Bert based Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.83 0.80
+StructBERT-mental (trained by swdd origin) 0.81 0.84 0.69 0.76 0.85 0.86
+StructBERT-mental (trained by wu3d) 0.85 0.85 0.92 0.92 0.83 0.80
+Llama3-8b 0.86 0.86 0.79 0.79 0.60 0.63
+Glm4-9b-chat 0.76 0.81 0.77 0.81 0.79 0.82
+Qwen2.5 7b 0.88 0.89 0.84 0.85 0.94 0.94
+GPT-4o-Mini 0.65 0.74 0.65 0.74 0.53 0.68
+GPT-4o 0.92 0.92 0.91 0.92 0.83 0.86
+DeepSeek-R1 671b 0.86 0.87 0.89 0.90 0.81 0.84
+Table 7: Model Classification Performance Comparison on Different Test Datasets.
+A.7.  Module II Prompt
 Figure 4: Module II Prompt.
-A.6 Prompt for Table 3 On Section 5.3.1
-Prompt is as follows: You are a professional in psychology and text sentiment analysis. Please determine whether the user is at risk of depression based on the following six dimensions:
-  • Negative emotions (secondary criterion)
-  • Depressive psychological state (primary criterion)
-  • Depression-related clinical symptoms (primary criterion)
-  • Potential external factors causing depression (secondary criterion)
-  • Depression-related medical expressions (primary criterion)
-  • Depression-related language expression patterns (secondary criterion)
-For each dimension, please cite the corresponding text reference and provide a concise explanation.
-[Output Data Example]
-  • Answer: “Yes” or “No.”
-  • Briefly explain the manifestations in each of the six dimensions, and cite the corresponding text reference.
-  • In conclusion, “The user is at risk of depression” or “The user is not at risk of depression.”
-A.7 Bert type models
-BERT-base-chinese [^9] https://huggingface.co/google-BERT/BERT-base-chinese is built on the BERT-Base architecture Devlin et al.
-(2019) and is pre-trained on extensive Chinese corpora.
-It adopts masked language modeling (MLM) and next sentence prediction (NSP) to learn deep bidirectional representations of Chinese text at the character level.
-Chinese-RoBERTa-wwm-ext Cui et al.
-(2019) leverages RoBERTa Liu et al.
-(2019) with a whole word masking (WWM) approach.
-Instead of masking individual characters, it masks entire words during MLM, leading to a richer understanding of context.
-The extended pre-training (ext) on large Chinese corpora further refines its effectiveness in various NLP tasks.
-StructBERT-mental Wang et al.
-(2019) is based on StructBERT, which incorporates linguistic structures into its pre-training.
-By focusing on word and sentence structure, and adding specific pre-training for mental health texts, it improves performance in identifying and categorizing mental health-related content.